### v1

# OODTE: A Differential Testing Engine for the ONNX Optimizer

Abstract
With $700$ stars on GitHub and part of the official ONNX repository, the ONNX Optimizer consists of the standard method to apply graph-based optimizations on ONNX models. However, its ability to preserve model accuracy across optimizations, has not been rigorously explored. We propose OODTE, a utility to automatically and thoroughly assess the correctness of the ONNX Optimizer. OODTE follows a simple, yet effective differential testing and evaluation approach that can be easily adopted to other compiler optimizers. In particular, OODTE utilizes a number of ONNX models, then optimizes them and executes both the original and the optimized variants across a user-defined set of inputs, while automatically logging any issues with the optimization process. Finally, for successfully optimized models, OODTE compares the results, and, if any accuracy deviations are observed, it iteratively repeats the process for each pass of the ONNX Optimizer, to localize the root cause of the differences observed. Using OODTE, we sourced well-known $130$ models from the official ONNX Model Hub, used for a wide variety of tasks (classification, object detection, semantic segmentation, text summarization, question and answering, sentiment analysis) from the official ONNX model hub. We observed $9.2$% of all model instances presenting issues leading into the crash of the optimizer, or the generation of an invalid model while using the primary optimizer strategies. In addition, $30$% of the classification models presented accuracy differences across the original and the optimized model variants, while $16.6$% of semantic segmentation and object detection models are also affected, at least to a limited extent. Finally, models related with text analysis and generation were found to be generally robust to the optimization process. Overall, we detected 15 issues, 14 of which were previously unknown, associated with optimizer crashes and accuracy deviations observed in $9$ out of the $47$ available optimizer passes but also the optimizer as a utility in general, which we reported to the developers of the ONNX Optimizer. Our method is simple, yet effective as indicated by the results, and can be easily adapted to other AI model optimization tools and compilers, even beyond the ONNX domain.
I Introduction
Artificial Intelligence has led to a technological revolution, not only within Computer Science but across various aspects of daily life. Deep Learning (DL), primarily implemented through Deep Neural Networks (DNNs), is applied across a wide range of domains, including autonomous vehicles and medical imaging. As these systems become more prevalent and are assigned increasingly complex tasks, the demand for their computational efficiency intensifies. To address this, both researchers and practitioners have developed tools and environments aimed at optimizing DL models [13, 17]. Additionally, widely-used AI compilers and frameworks, such as Apache TVM [12] and MLIR [32], incorporate their own native sets of optimization passes. These include standard techniques such as operator fusion and constant folding, as well as more aggressive approaches like reduced-precision arithmetic, including the use of fast-math optimizations, which may compromise numerical accuracy.
Models can also be represented using industry-standard formats, such as ONNX [9]. By applying transformations to models in these formats, it is possible to reduce model size and improve performance. For ONNX in particular, the ONNX Optimizer [5] serves as the default tool for performing graph-level optimizations. As part of the official ONNX toolkit, it focuses on simplifying model structure while aiming to preserve functional correctness, particularly in terms of accuracy. However, the degree to which these optimizations maintain model accuracy remains an open question. The ONNX ecosystem also includes additional tools that can further optimize models, potentially at the cost of accuracy—such as the ONNX Quantizer [2]. Developers are typically aware of these trade-offs and apply such tools with consideration of their impact.
We investigate the extent to which ONNX Optimizer transformations affect the functional correctness of DNN models. Ideally, such transformations should not introduce any deviations in model behavior. To this end, we propose the OONX Optimizer Differential Testing Engine (OODTE), a framework for performing differential analysis between original and optimized models. Using OODTE, we collected 130 models of various types, including classification, object detection, and text generation from the official ONNX Model Hub [8]. Each model was optimized using ONNX Optimizer and evaluated on standard benchmark datasets to compare accuracy before and after optimization. In cases where discrepancies were observed, we conducted pass-by-pass optimizations to isolate the specific transformations responsible for the faults.
We focus on the ONNX Optimizer due to its impact and popularity (approximately $700$ GitHub stars) as a model optimization tool. In addition, it is integrated into the main ONNX repository and serves as a foundation for other widely adopted utilities, such as the ONNX Simplifier [6], which has received over 4,000 GitHub stars. However, existing ONNX Optimizer tests primarily focus on toy models rather than real-world models, primarily used for unit testing purposes. We aim to address this gap by introducing a differential testing methodology through OODTE. However, our approach is not limited to ONNX; it is generalizable to other AI compilers and optimization frameworks. Our methodology builds upon established principles of differential testing [45], which have traditionally been employed for detecting inconsistencies across software variants or transformations.
In total, we observed accuracy deviations in $34$ out of the $130$ models ($26.1$%). Additionally, we encountered issues with $12$ model instances ($9.2$%) while using the primary (fuse and eliminate) optimizer passes - with either the optimizer crashing or generating an invalid model. Through detailed analysis, we identified $15$ distinct bugs, of which only one had been previously reported—and even then, only partially and without a conclusive explanation. In that case, we were able to provide substantial additional information not previously documented. All identified bugs were reported to the official ONNX Optimizer repository, accompanied by detailed descriptions, so that the optimizer developers can further examine and resolve them.
Overall, we present the following contributions:
1. OODTE  - a differential testing tool to examine the effects of the ONNX Optimizer to DNNs.
2. A survey of 130 models optimized using the ONNX Optimizer, presenting the effects of the optimizations on model correctness in detail.
3. A set of faults detected using OODTE - reported to the official ONNX Optimizer repository.
II Background
II-A Deep Learning Models
Deep Learning Models (DNNs) are a subset of AI models widely used for prediction tasks across various domains. We focus on the following types of models:
II-A1 Image Classification
One of the earliest and most prevalent applications of DNNs is image classification, typically performed using Convolutional Neural Networks (CNNs). These models employ convolution operations, which function as filters to extract specific features from the input image, using weights and biases, of which values are formed during training. Image classification models are often of small-to-medium size in terms of parameters and layers, taking an image (often preprocessed) as input and outputting a list of top-K prediction pairs (prediction, certainty_score). The sum of the $K$ certainty_score instances equals $1$, with the final operation commonly being a Softmax [47] to amplify the highest prediction probability.
II-A2 Object Detection & Semantic Segmentation
Object detection models extend image classification models by not only classifying objects but also detecting where they are placed within an image, typically using bounding boxes. These models also heavily rely on convolution operations. Object detection is widely used in domains like autonomous driving, where multiple objects need to be detected within a scene. A subset of object detection tasks is semantic segmentation, which classifies each pixel in an image into regions corresponding to different object categories. Each pixel is assigned a specific semantic label. For example, a semantic segmentation model for an autonomous vehicle might label parts of a scene as ”road,” ”cars,” ”pedestrians,” etc., providing a more detailed understanding of the environment than simple object detection.
II-A3 Text Generation
Text generation models are a class of models focused on tasks such as text recognition, summarization, sentiment analysis, and more. These models function similarly to image recognition models, receiving an input and generating an output prediction. In text generation, the model typically receives a sequence of natural language input tokens, generates a prediction for the next token, appends it to the input, and discards the leftmost token of the sequence. This process continues iteratively until a stopping condition is met. This approach is referred to as the sliding window technique.
Text generation models are a subset of Generative AI models, responsible for generating content. A relatively new but rapidly evolving family of models in this category are transformers [54], which use an architecture that processes input data to create internal representations, capturing context. These representations are then used to generate new information, typically in the form of text, by predicting one token at a time sequentially (using the sliding window technique described earlier) until the model output is fully generated.
II-B Graph Optimizations
For deep neural networks (DNNs) to perform effectively on complex tasks, they must be optimized to operate more efficiently while maintaining their performance. A common approach to achieving this is optimizing the model graph to improve computational efficiency. This may involve techniques such as node pruning, node fusion, and constant folding, among others. Similar to conventional software, these optimizations are applied to DNNs, with AI compilers [12, 32] incorporating such techniques. Additionally, tools like the ONNX Optimizer, which we examine in this work, are specifically designed for this purpose.
II-C Comparison Methods
Based on the model type under test, OODTE utilizes a set of well-known and established comparison methods. In particular:
  • For classification models, we use Kendall’s Tau Rank Correlation [28], which measures the correlation between ranked lists. This method is suitable for classification tasks, as it compares the order of ranks, yielding a correlation score in the range of $-1$ (completely negative correlation) to $1$ (completely positive correlation). Kendall’s Tau penalizes order discrepancies between the ranks. In classification tasks, each model typically generates a list of top-K predictions per input. We can then extract these top-K predictions from two model variants and apply Kendall’s Tau to compare their outputs.
  • For Object Detection and Semantic Segmentation models, we utilize the following well-known metrics: (1) Average Intersection-over-Union (IoU) – measures the overlap between predicted and ground truth regions, indicating how well a model recognizes class labels. (2) F1 Score – the harmonic mean of precision and recall, balancing false positives and false negatives. (3) Average Precision (AP) – evaluates how many of the model’s positive predictions were actually correct, focusing on precision across different thresholds. (4) Average Recall (AR) – measures how many actual positives the model correctly identified, focusing on recall across thresholds. (5) Mean Average Precision (mAP) – aggregates the model’s performance across multiple thresholds, evaluating both detection and localization accuracy. To account for varying levels of detection confidence, we performed calculations for metrics (2)-(5) using three different IoU thresholds: $0.5$, $0.75$, and $0.9$. These thresholds are necessary to assess model performance at different levels of localization accuracy, where higher thresholds (e.g., $0.9$) require more precise object detection in the image.
  • For text generation models, we utilize BLEU Score [46], which measures the precision of n-gram matches between the generated text and reference text, evaluating the quality of the generated output. The base refers to the output from the original model, and each optimized model’s output is compared to this base to assess similarity.
  • For binary sentiment analysis, we compare the predictions of each sample between the original and optimized models, calculating the percentage difference in predictions to assess the impact of optimization on sentiment classification.
III System Architecture & Methodology
Figure 1: Architecture of OODTE  with the following primary modules: (1) Model Orchestrator Module, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module.
The architecture of OODTE  with its different modules is shown in Figure 1. OODTE  has the following intuitive workflow: load models, optimize them, execute both the original and the optimized variants across the dataset and compare the results. If any discrepancies in model outputs are detected, OODTE  performs an iterative approach where each model under test is optimized with each ONNX Optimizer pass in isolation, and then the results are compared with the original model. This is done to localize the optimization pass responsible for causing an output difference. All results are stored in report files, both for reporting and localization purposes. In detail, OODTE  consists of the following primary modules: (1) Model Orchestrator, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module. OODTE  loads model configuration settings in two ways: (1) via local configuration file, and (2) by fetching configuration via the ONNX repository from which the models are loaded from. This is managed by the Model Orchestrator Module. The modules within OODTE  in Figure 1 is explained in the following Sections.
III-A Model Orchestrator Module
The Model Orchestrator Module is responsible for all tasks related to fetching, loading (including model configuration) and preparing models in order to get optimized and run. The module supports two loading options for DNN models: (1) local loading of ONNX models, and (2) loading of an online repository. At the moment, OODTE supports the official ONNX Model Hub [8], however, as it utilizes the official Hub API, it can trivially support third party ONNX model repositories. In addition, the module supports fetching and loading of multiple models simultaneously, while it also performs model caching when fetching from a model hub, as well as checksum verification. Once each model and its respective configuration are properly loaded, they are propagated to the Optimizer Module and the Runner Module.
III-B Optimizer Module
The Optimizer module is responsible for the generation of optimized models, using the ONNX Optimizer at its heart. It supports two settings: (1) full optimization, applying all the ONNX Optimizer passes to the model under test, and (2) partial optimization, where it only applies a specific set of passes to the model. It is worth noting that in (1), by optimizer design, only the fuse and eliminate passes are applied, leaving out a number of secondary passes (e.g., ‘rename_input_output‘). OODTE supports the utilization of these passes for testing, in order to explore all optimizer operations. This is useful across OODTE pipeline, when faulty behavior is discovered across model evaluation and the models presenting the faults need to be examined in depth. OODTE isolates such cases after a full pass optimization exeuction and performs per-pass optimization generation for that purpose.
Once a model is optimized with either setting, it then gets propagated to the Runner model, in order to perform inference.
III-C Runner Module
The Runner model does precisely what its name suggests: executes the models under test. In particular, it receives both the original and the optimized model under test, as well as a dataset that will be used as input for both model variants. It then performs model inference, using ONNX Runtime [3], combined with the necessary model preprocessing, while utilizing the model configuration loaded in the Model Orchestrator component. OODTE supports different types of datasets, based on the type of models being examined - from images to IMDB reviews. It can load the inputs from local files, but also fetch the dataset from well-known packages, such as the datasets [4] package.
In order to be able to utilize large-size datasets and extract results from heavier models while running on low-end devices, the Runner Module supports execution in-parts: by definition in OODTE configuration, a run can be performed in $C$ chuncks, where each chunk consists of $dataset/C$, with $dataset\%C$ also being the final chunk of the run. This applies both for the original and the optimized model, and the results are extracted and propagated to the Metrics Comparison Module, for evaluation and reporting. Following these approach, we were able to conduct our survey on two devices designed for home use without relying on a cluster for results, described in Section V.
III-D Metrics Comparison Module
Once the run completes per-chunk for both the original and the optimized model, the results are propagated to the Metrics Comparison Module. This module performs comparison against the results, by selecting the appropriate comparator based on the model type. For classification types, OODTE supports first-label comparison, but also calculates the Kendall Rank Correlation Coefficient across the $top-K$ predictions - commonly known as Kendall’s Tau [28]. For Object Detection models, it utilizes the same methods of classification for labels and ranks, but also applies Intersection over Union (IoU [68]), Precision, Recall, and F1 Score [15] given an IoU threshold. For text generation models (e.g., GPT-2), it utilizes BLEU [46] score in order to compare the text generated out both the original and the optimized model, as well as direct tensor comparison for cases of question and answering (Q&A). In addition OODTE is build in a manner that new comparators can be added, in cases that a new category of models is tested.
Once a method is selected, the Metrics Comparison Module calculates the metrics based on model type and generates a report with all results, in JSON format. The results can then be visualized, but also they are utilized from OODTE in order to localize faults. As aforementioned, if a difference is detected across the original and the optimized model, then OODTE retrofits the reported data to the Optimizer Module. Then, the Optimizer Module generates model optimizations per-pass and repeats the process, reporting how each model behaves when optimized on each pass in isolation. These results are also stored in separate reports. The whole process is iterative for all chunks of input data, for each optimization (if an optimized model is found to be problematic) and for all models under test.
IV Related Work
Conventional compiler testing is a domain that has been explored [10, 74, 16], involving differential testing [49, 67], as well as including compiler optimizations [11, 27, 19]. However, in terms of testing AI tools and compilers, most of the existing research has focused on (1) the direction of compiler fuzzing and (2) examining context-specific AI tools and APIs, such as DL framework converters.
In terms of compiler fuzzing, HirGen [43] is a utility for Deep Learning compiler fuzzing, focusing on the high-level model intermediate representation (IR) of TVM. TVMFuzz [1] allows generic-purpose TIR fuzzing on Apache TVM, detecting primarily crash and bit manipulation bugs. Tzer [38] is a coverage-guided fuzzer that applies modifications on models compiled in Apache TVM utilizing low-level IR (Tensor IR). GenCoG [70] focuses on generating valid and expressive TVM computation graphs by defining a domain specific language for constraint definition, as well as concolic constraint solving. NNSmith [37] focuses on valid DNN graph generation, usable to examine the compiler for bugs similarly with traditional grammar-based fuzzing techniques (e.g., [74, 16]), and MLIRSmith [69] integrates NNSmith’s fuzzing capabilities in MLIR [32] randomly. Suo et al. [62] also follow an operation dependency approach to guide the fuzzing process in MLIR. DeepDiffer [35] applies low-level IR fuzzing in a coverage-guided manner and comparing results across different DL compilers. SYNTHFUZZ [34] combines grammar-based fuzzing with custom mutation synthesis in order to effectively fuzz MLIR dialects. Li et al. [33] propose a methodology to uncover hardware-related faults by generating fault-sensitive, adversarial tests. Finally, Su et al. [61] fuzz DL compilers in the context of dynamic features, detecting bugs in PyTorch and its underlying tensor compiler.
Regarding testing techniques for context-specific AI tools and APIs, CRADLE [48] performs an execution analysis on the model graph in order to detect faults, while LEMON [71] utilizes the metrics used by CRADLE and applies mutation testing in order to identify issues in the model. FAME [59] is based on LEMON but enables API mutation and optimizations for layer and weight mutation, guided by the detection of inconcistency faults. Similarly, Audee [20] aims to detect logical, not-a-number bugs, and crashes by applying an exploratory approach in combination with mutation testing. Finally, Louloudakis et al. propose FetaFix [42] - a utility that allows automatic fault detection in model converters, involving ONNX as an intermediate model representation.
Overall, limited work has been done in the direction of testing AI compiler optimizations. Choi et al. [14] have conducted an empirical study towards optimizer comparisons, however they do not consider graph-based optimizations in the post-training phase such as the ONNX Optimizer) as they focus on stochastic optimizers focusing primarily on the model training process, such as Adam [29]. DiffChaser [73] explores the application of differential testing across optimized and unoptimized DNN model variants. Louloudakis et al. propose DeltaNN [41], a utility to explore model robustness across different configurations - including but not limited to compiler optimizations. Xiao et al. [72] propose a metamorphic testing methodology for AI compiler bug detection, including issues related to optimizations.
To the best of our knowledge, OODTE is the first that thoroughly tests the ONNX Optimizer (and AI optimizers in general), using differential testing in order to detect faults. The closest existing work to our methodology, is DiffChaser [73], DeltaNN [41] and FetaFix [42]. However, these tools considerably differentiate from our contribution. In particular, DiffChaser attempts to detect problems related to AI optimization techniques (such as quantization) related to the models under test, not exploring potential faults introduced from bugs in the optimizer. DeltaNN applies differential testing, but focuses on uncovering behavioral inconsistencies in the models under test when deployed using different configurations, rather than focusing on compiler issues. FetaFix, on the other hand, aims to uncover issues related to converters using differential testing and involving ONNX. Its approach however does not consider model optimizations or AI optimizer tools.
V Experiment Setup
For our experiment set, we sourced $130$ model instances from the ONNX Model Hub. In particular, we selected $93$ models utilized for classification tasks, $30$ models used for object detection and semantic segmentation tasks, and $7$ transformer models [54] utilized for text generation tasks, such as question and answering, (BERT-SQuAD [18]), sentiment analysis (RoBERTa [40]), text summarization (T5 [51]) and generic-purpose text generation (GPT-2 [50]). We utilize models of a variety of opsets based on hub availability, ranging from $7$ (the lowest compatible with the latest ONNX Runtime version by the time of this publication - ONNX Runtime v1.21.11.21.11.21.11.21.1) up to $12$ (the highest provided from ONNX Model Hub via their API). However, our approach is generalizable and not restricted to one hub or limited by opset versioning. We selected it because it consists of a standard set of models, refined and prepared for wide use by developers and researchers. The same applies for the models we selected: we focused on models that are mature and strong representatives of their categories (e.g., MobileNet [24], EfficientNet-Lite4 [65] DenseNet121 [25], InceptionV2 [63] RetinaNet [36], Mask/Faster R-CNN R-50-FPN [21, 56] SSD [39], RoBERTa [40] and GPT-2 [50]), but our methodology can apply to any type of ONNX model. In addition, we utilized text generation models that generated results in a direct, one-off manner (such as RoBERTa), but also in a token retrofitting manner (such as GPT-2). In the latter case, we considered selection of the first, second, and third token prediction, so that our results are as comprehensive as possible.
In terms of computational resources, we ran our models on mid-range systems with Intel i5 and i7 CPUs, reflecting the typical hardware used by developers during model development and fine-tuning. Our method is generalizable to larger models, limited only by ONNX model size constraints, and can be extended to high-end systems with a trivial extension for hardware acceleration (using the ONNX Runtime API [3], or an AI compiler such as Apache TVM [12]).
We selected datasets appropriate to each model task. For classification and object detection, we used the ILSVRC2017 [57] validation dataset, containing $5500$ images, commonly used for benchmarking image recognition models. For text generation, we used SQuAD-v2 [52], which includes $11873$ questions and contexts, to evaluate Q&A tasks and general-purpose transformer behavior. Additionally, for summarization and sentiment analysis, we used the IMDB dataset [44, 66], containing $25000$ entries.
To accurately detect differences across model types, we employed comparison methods appropriate for each task. Specifically, we used the following approaches:
  1. 1.
    For image recognition models performing classification tasks, we utilized Kendall Rank Correlation Coefficient [28] (KT), measuring differences for top-K prediction labels of each dataset input image, for $K=1,5$, and $10$. If the KT value was found to be $<1$ for each input, we counted this input as different.
  2. 2.
    For image recognition models performing object detection and semantic segmentation, we calculated precision, recall, and F1 score for each dataset input, and extracted their average, combined with different Intersection-over-Union (IoU) thresholds ($IoU=0.5,0.75$, and $0.9$), as well as IoU, mean Average Precision (mAP), along with average precision and recall. We considered the original model output as the base, and the optimizer model the one under test for our metric calculations, and the higher the values observed, the better the optimized model aligns with the original.
  3. 3.
    For text generation models performing question answering and summarization, we utilized BLEU score [46] across output texts for each input, monitoring instances where $BLEU\_Score<1$ for each input.
  4. 4.
    For sentiment analysis, we compared each output value consisting of ${0,1}$ across the original and the optimized variants, registering any differences observed.
Finally, we evaluated all existing ONNX Optimizer passes. However, OODTE can also be applied to new or custom passes. Moreover, our methodology can be extended to other optimizers or model optimization APIs (e.g., Apache TVM), provided that the API supports applying passes both individually or as a bundle.
VI Results
Following the execution of the full set of experiments, OODTE successfully identified 15 bugs. Of these, 14 were entirely new (not previously reported on the ONNX Optimizer issue tracker), while one was a previously reported issue that lacked sufficient detail, including model information and reproducibility steps. We provided comprehensive information and full reproducibility details for this issue.
The issues were associated with $9$ of the $47$ passes in the optimizer. However, not all passes are of equal importance; by default, the optimizer utilizes passes related to fuse and eliminate operations. Therefore, $8$ of the discovered issues are deemed more critical, as they are part of the optimizer’s core bundle (default) strategy. Of these, $6$ are directly linked to specific passes (e.g., fuse_bn_to_conv, which fuses batch normalization nodes with convolutions), while the remaining issues were general and related to the optimizer’s overall usage. We present the impact of these faults both on the optimization process and on the correctness of the optimized models.
TABLE I: Issues detected using OODTE, along with models manifested at. Opset version indicated near each model name.
# Pass Error Type Models
1 lift_lexical_references Removes used value references. Crash EfficientNet-Lite4 (11), SSD (12)
2 split_init List index out of range Crash YOLOv3 (10, 12)
3 split_init split_predict Field is required but missing. Crash DenseNet121 (7-9), YOLOv2 (10), SSD (12), Inception v1 & v2 (7-9), Tiny-YOLOv3 (11) YOLOv3 (10, 12)
4 split_predict Required inputs are missing. Crash SSD (12), DenseNet121 (7), InceptionV2 (6-9)
5 split_predict Bad spec for node. Crash YOLOv3 (10), DenseNet121 (7) InceptionV2 (6-9)
6 split_predict Input element number mismatch. Crash FCN ResNet-50-int8 (12)
7 eliminate_shape_gather Indices value not equal to node dims size Crash Faster R-CNN R-50-FPN (10)
8 All Passes Models present slightly different results. Diff. See Table II
9 rewrite_input_dtype Output name used multiple times. Crash GPT-2 (10)
10 eliminate_ consecutive_ idempotent_ops Incompatible Shape Error. Crash (Run) Mask R-CNN/Faster R-CNN: R-50-FPN-int8 (12), R-50-FPN-qdq (12), R-50-FPN-fp32 (12)
11 eliminate_ duplicate_ initializer Erase output assertion failed. Crash Faster R-CNN R-50-FPN-qdq (12)
12 fuse_bn_into_conv Duplicate initializer introduced. Crash (Run) RetinaNet (9)
13 eliminate_shape_gather eliminate_nop_reshape Added initializers not used by the model. Warning Faster R-CNN R-50-FPN-int8 (12)
14 All Passes Invalid input scale. Crash VGG-16-BN (7), VGG-19-BN (7), ResNet101_DUC_HDC (7)
15 fuse_bn_into_conv Bad graph structure Diff. See Table II
VI-A Optimizer Crashes
Out of the $15$ issues, $12$ were related to model crashes, $2$ caused model accuracy deviations and $1$ obfuscated the model with redundant initializers, leading to a runtime warning. In particular, we observed that those crashes concerned $8$ out of the $47$ passes ($17$%), and occurred for specific models, with object detection models being more prone to errors. In particular, R-CNN models, such as Faster R-CNN R-50-FPN [56], RetinaNet [36], VGG-16, VGG-19 [60], as well as YOLOv3 [55], presented crashes by using the optimizer, in a range of opset instances and in passes such related to the modification of gather nodes, as well as the elimination of idempotent operations. In total, we observed $31$ model instances that either the optimizer crashed upon processing them, or the optimized model was problematic, leading into a runtime crash. However, some secondary passes are known to be unstable (e.g., split_init and split_predict describe in their documentation that they not strictly comply with the ONNX format), and therefore we do not focus on them as sources of faults in the optimization process. As a result, from those, $12$ model instances ($9.2$%) while using the primary (fuse and eliminate) ONNX Optimizer passes and are of higer concern.
Overall, classification models were affected towards optimizer crashes to a much lesser extent. However, they were considerably affected in terms of correctness by a subset of faults that did not lead to a crash - which we present below. From text generation models, one issue was related to the crash of the optimizer, and it concerned the rewriting of the input type. However, other text generation models did not present any crash behaviors while optimized using the ONNX optimizer.
VI-B Deviations In Optimized Model Accuracy
Out of the $130$ model variants tested, we observed unexpected accuracy differences between the original and optimized models. This was an unexpected outcome, since ONNX Optimizer transformations are graph-based and do not alter numeric precision, meaning no accuracy deviations were anticipated. Specifically, of the $93$ classification models, $28$ showed accuracy differences, while $5$ of $30$ object detection models had discrepancies. Only one transformer instance (GPT-2) exhibited negligible differences ($<0.001\%$), with all others showing identical results. The differences primarily affected the top-5 and top-10 predictions, with no changes in the top-1 predictions, except for the SSD model, which showed a $2.9\%$ difference in top-1 prediction. Although top-1 accuracy is often prioritized, it is important to recognize that changes in lower-ranked predictions can also indicate issues. Our methodology identified these inconsistencies, which have been reported to the optimizer developers. For instance, ResNet50-caffe2 showed a $12.9\%$ difference in top-10 labels and $2\%$ in top-5 labels, while ShuffleNet presented $0.5\%$ and $1\%$ differences in top-5 and top-10 predictions, respectively.
Figure 2: Illustrative example of problematic label classification and object detection between original (blue) and optimized (orange) model variants, with the bounding box being affected in both cases, with the same (left) and different (right) object label prediction, along with their affected certainty scores.
Figure 3: Model graph malformation caused by the fuse_bn_to_conv pass in EfficientNet-Lite-4.
TABLE II: Observed differences across classification predictions for classification (CL) and object detection (OD) models.
Model Type Opset(s) Top 1 Top 5 Top 10 Issue
AlexNet [30] CL 7-9 0% 0% 0.02% Format Version Change
DenseNet121 [25] CL 7-9 0% 0% 0.02% Format Version Change
EfficientNet-Lite4 [65] CL 11 0% 0.04% 1% Graph Malformation
GoogleNet [64] CL 7-9 0% 0% 0.04% Format Version Change
Inception-1 [63] CL 7-9 0% 0% 0.25% Format Version Change
Inception-2 [63] CL 7-9 0% 0.02% 0.05% Format Version Change
MobileNet v2 [24, 58] CL 7-9 0% 0% 0.02% Format Version Change
ResNet18 [22] CL 7 0% 0.02% 0.02% Format Version Change
ResNet34-v2 [23] CL 7 0% 0.02% 0.02% Format Version Change
ResNet50-caffe2 [22] CL 7-9 0% 2% 13% Format Version Change
ShuffleNet-v1 [75] CL 7-9 0% 0.7% 1% Format Version Change
SqueezeNet 1.0 [26] CL 7-9 0% 0% 0.02% Format Version Change
SSD [39] OD 12 2.9% 4.3% 6% Graph Malformation
Following runs with OODTE, we concluded that $2$ primary issues were related with accuracy deviations across models:
  • The optimizer automatically upgrades ONNX formats of $v3$ to $v4$, without any warning or indication. This affects models of opsets $7$, $8$, and $9$, which is an outdated opset, but still backwards compatible to the current ONNX Runtime version (v1.21.1), well-represented in the ONNX Model Hub, with $41\%$ of the models in it consisting of any of these opset versions. Of course, not all models within this opset range are affected, yet we found a considerable amount of models presenting differences across their original and optimized variants. We have reported this issue to the optimizer developers and they have acknowledged this behavior for further inspection.
  • The pass fuse_bn_to_conv essentially fuses BatchNormalization and Convolution operators within the model. We observed that this pass was found problematic in a number of cases, resulting in both a crashing bug and an issue where redundant nodes are introduced to the model. In the latter case, the pass attempts to fuse BatchNormalization and Conv nodes, but does so in an incorrect manner, essentially breaking both node properties into chunk nodes. The result can be visually observed in Figure 3. We found that a number of models were producing different accuracies when optimized using this pass. This issue was previously encountered by another user of the optimizer, however no details were given regarding how this issue could be reproduced or what models were affected by it. We raised a comprehensive issue to the ONNX Optimizer developers, providing all the necessary information and details related to its behavior.
Table II showcases the accuracy differences detected in top-K labels for classification and object detection moels, indicating which of the above issues is responsible for this behavior. As an example, Figure 2 illustrates how misclassification and bounding box issues manifest in an input image, with relation to the original (blue) and the optimized (orange) model.
Model Op. Avg. F1 Avg. Prec. Avg. Recall mAP Avg. IoU
SSD [39] 12 95.28% 95.28% 95.28% 86.62% 97.37%
Tiny YOLOv3 [55] 11 99.56% 99.56% 99.56% 90.57% 99.61%
YOLOv2 [53] 9 99.95% 99.95% 99.95% 90.85% 88.50%
YOLOv3 [55] 10 98.83% 98.83% 98.83% 86.66% 98.78%
YOLOv3 [55] 12 98.83% 98.83% 98.83% 86.65% 98.78%
(a) Object Detection Scores with IoU threshold of 50%.
Model Op. Avg. F1 Avg. Prec. Avg. Recall mAP Avg. IoU
SSD [39] 12 95.27% 95.27% 95.27% 86.6% 97.37%
Tiny YOLOv3 [55] 11 99.36% 99.36% 99.36% 89.97% 99.61%
YOLOv2 [53] 9 99.97% 99.97% 99.97% 90.87% 88.60%
YOLOv3 [55] 10 97.87% 97.87% 97.87% 85.71% 98.78%
YOLOv3 [55] 12 97.87% 97.87% 97.87% 85.71% 98.78%
(b) Object Detection Scores with IoU threshold of 75%.
Model Op. Avg. F1 Avg. Prec. Avg. Recall mAP Avg. IoU
SSD [39] 12 95.27% 95.27% 95.27% 86.6% 97.37%
Tiny YOLOv3 [55] 11 99.34% 99.34% 99.34% 89.97% 99.61%
YOLOv2 [53] 9 99.97% 99.97% 99.97% 90.87% 88.60%
YOLOv3 [55] 10 97.8% 97.8% 97.8% 85.69% 98.78%
YOLOv3 [55] 12 97.8% 97.8% 97.8% 85.69% 98.78%
(c) Object Detection Scores with IoU threshold of 90%.
VI-C Optimizer Warnings
We observed two issues related to the obfuscation of the optimized models with unused initializer entries, associated with the eliminate_shape_gather and eliminate_nop_reshape passes. When executed with ONNX Runtime, these entries were ignored, triggering a warning in the format: Removing initializer ’INITIALIZER_ID’. It is not used by any node and should be removed from the model. Upon manual inspection, we found no significant differences in model size. However, this issue warrants further investigation for potential side effects. We have reported it to the optimizer developers for additional review.
VII Tool Usability
OODTE  can be used directly to test whether the ONNX Optimizer generates optimized models without negatively affecting their accuracy. Specifically, OODTE can verify that the ONNX Optimizer performs correctly on real-scale models of various types and architectures. Currently, OODTE supports differential testing for classification, object detection, and text generation models. However, OODTE can also be easily extended to support new model types, as long as (1) the models can be compiled into the ONNX format for optimization and (2) a comparator is provided to compare the results between the base and optimized models for the new model type.
OODTE is an automated utility designed to test the ONNX Optimizer. It can be configured to perform tests both manually and automatically. As a configuration-based tool, OODTE can fetch models and datasets automatically, making it adaptable for use in a Continuous Integration pipeline. This allows OODTE to test the ONNX Optimizer whenever new functionality (e.g., new passes) is added, whether by the optimizer developers or users implementing custom passes.
In terms of models and datasets, OODTE is adaptable to models provided both locally and fetched from online repositories. By default, OODTE loads models from the official ONNX Model Hub [8], which was used in our experiments. However, by configuring OODTE to use a custom ONNX Hub that follows the ONNX API specification for model repositories [7], it can fetch and utilize models without further modification. The same flexibility applies to datasets. OODTE can fetch datasets using widely used packages like the datasets [4] package or from local storage. Additionally, OODTE can be easily extended to support any dataset suitable for testing.
Finally, while OODTE applies a differential testing methodology focused on the ONNX Optimizer, this approach is generalizable and can be easily adapted to other AI optimizers or compilers that meet two primary requirements: (1) they provide a runtime for model execution, and (2) they expose an API that includes all the passes applied during model optimization, both as a bundle and in isolation.
VIII Threats To Validity
In terms of threats to validity, we considered and addressed the following aspects: (1) Comprehension of Experiment Set: We sourced $130$ models from the official ONNX Model Hub, covering classification, object detection, and text generation—categories widely used across various domains. These models are well-established and mature, as the main purpose of the hub is to provide researchers and developers with easily accessible models that require minimal configuration and fine-tuning. The hub offers a broad range of ONNX model versions (opsets), which we considered in our experiments, ranging from version $7$ (the lowest version compatible with the current ONNX Runtime at the time of writing) to version $12$ (the highest available via the ONNX Hub API).
We believe our experiment set is a representative and comprehensive set of models for testing and experimentation. However, OODTE supports models from any hub that complies with ONNX API specification. While our experiment set does not include all models in the hub, we believe it is representative. We excluded models with an opset version lower than $7$ (which are incompatible with the current ONNX and ONNX Runtime version), as well as models for body analysis and super-resolution, since the selected classification and object detection models adequately cover these areas. Additionally, models that only prepare inputs for preprocessing were excluded for being non-generalizable, as were certain transformer models, which were represented by GPT-2.
Overall, we believe our experiment set is comprehensive, representing models commonly used by developers and researchers in the community and industry across a wide range of domains, while adhering to the latest versions supported by ONNX. (2) Optimization Order: We acknowledge that the optimization order can impact model performance [31], a common challenge in compiler research. However, exploring this aspect is beyond the scope of this work. To ensure the validity of our approach, we used two modes in OODTE: (1) the default ONNX Optimizer configuration (which includes the fuse and eliminate passes), and (2) runs where all $47$ optimizer passes were applied atomically. While our approach does not address potential defects arising from the combination of passes, it has been shown to be effective towards detecting correctness issues. (3) Proper Optimization Usage: We followed all guidelines provided by the ONNX Optimizer API to ensure the correct use of each pass. In cases where a pass was expected to produce model graphs that were not strictly compliant with the ONNX format, we took note of this and clearly indicated such instances when reporting the respective issues to the developers. (4) CPU Devices Used: We used both an Intel i5 and an Intel i7 in parallel for our experiments. To ensure consistency across devices, we (a) randomly sampled a subset of models of each type and ran them on both devices, verifying that the results were consistent, and (b) referred to the work of Louloudakis et al. [41], who observed that accuracy remains consistent across different processing devices using optimization. This supported our finding from the sample that results remain consistent across devices.
IX Discussion
We make the following observations from our experiments:
(1) object detection models are more prone to lead into a crash during the optimization process. The vast majority of the crashes we detected were related to object detection and semantic segmentation models. We attribute this to the increased complexity of object detection models in comparison to those utilized only for classification tasks, which are typically simpler in terms of architecture.
(2) The object detection capabilities of problematic models might be considerably affected. In particular:
  • YOLOv3 models show a high average IoU, F1 score, precision, and recall (over 98%). The optimized model misplaces bounding boxes, but not enough to cause misclassifications, as no misclassifications were found in the top-10 label predictions.
  • YOLOv2 showed low IoU values, likely due to its use of multiple bounding boxes per detected object, where small deviations accumulate, negatively affecting the IoU score. However, the high F1 score, average precision, and recall suggest that these deviations do not significantly impact object detection, similarly to how it is illustrated by the left bird instance in Figure 2.
  • In contrast, SSD showed about $4.8$% deviation in F1, precision, and recall between the original and optimized models, leading to $2.9-6$% misclassification rates, despite a high IoU ($>97$%). This is similar to the right bird instance in the illustrative example in Figure 2.
All models presented a considerable mAP score deviation of up to $14.31$%. This indicates that some classes in the model were not properly identified, meaning that the optimizer might affect all object detection and semantic segmentation models towards missing specific objects. This can have a severe effect in models, as despite correctly classifying the objects they detect, they might ”miss” a non-negligible amount objects in the scenes presented to them.
(3) Classification models are more prone to differences in accuracy of the optimized model, compared to the original model. We observed two issues that had an adverse effect towards model classification accuracy. The root causes of such deviations, were attributed to two main issues, (a) one related to implicit, improper change of the ONNX format version for older model opset versions - in which case a large amount of classification models were built using these opsets, and (b) one related to the generation of problematic graph nodes, in relation to convolutional layers (fuse_bn_into_conv pass), a pass that has greater effect in classification models.
(4) Text generation models were affected only to a negligible extent. From the $8$ model instances under test related to text generation, we only found one crash issue related to GPT-2, as well as a very small number of samples (8) producing a BLEU score that indicates mismatch between the original and the optimized model, for the top-3 token selections of the model ($4$ when selecting the first token, $2$ when selecting the second and another $2$ when selecting the third token on each prediction). However, we consider this result negligible, as a common practice is to use random token selection across token predictions, and such a difference is so negligible it will essentially not affect model behavior.
Overall, we recognize that the ONNX Optimizer is a tool that performs well for most of the passes and produces good results. We aspire to contribute with this publication, so that the reported issues get fixed by the optimizer developers.
X Future Work
We identify three potential extensions of this work as future research directions:
(1) Execution Time Measurements: This study primarily focused on optimizer correctness, specifically detecting crashes and accuracy faults. However, given the context of model optimization, it is important to also assess how the optimizer impacts model performance, particularly in terms of execution time (for both CPU and GPU deployments) and other factors such as model size. OODTE can be easily extended to support hardware acceleration model execution (either by using the ONNX Runtime API [3], or by integrating a third-party AI compiler that generates target device code, such as Apache TVM [12]) We leave this as a subject for future work, which could provide valuable insights.
(2) Selection of Different Model Types: This work focused on classification, object detection, and text generation models. However, OODTE can be applied to other model types, such as speech recognition, super-resolution, and body recognition. We consider this as a direction for future work.
(3) Examine Optimization Pass Order: OODTE supports optimization with a subset of passes in a defined order. However, our current approach does not account for this, and it can be explored in future contributions.
XI Conclusion
We propose OODTE, a tool for the differential testing of the ONNX Optimizer. Using OODTE, we evaluated $130$ models across classification, object detection, and text generation tasks. We detected $9.2$% of model instances that crashed when using the primary optimizer strategies, while $30$% of classification models and $16.6$% of object detection models showed differences between the source and optimized versions, while text generation models remained robust. Additionally, OODTE identified $15$ previously unreported bugs related to crashes and accuracy degradation across $9$ optimizer passes, which we reported to the ONNX Optimizer developers.
XII Availability
For more information about OODTE, please contact the authors. We intend to release the tool’s source code in the near future.