### v2

# OODTE: A Differential Testing Engine for the ONNX Optimizer

Abstract
With over 700 stars on GitHub and being part of the official ONNX repository, the ONNX Optimizer is the default tool for applying graph-based optimizations to ONNX models. Despite its widespread use, its ability to maintain model accuracy during optimization has not been thoroughly investigated. In this work, we present OODTE, a utility designed to automatically and comprehensively evaluate the correctness of the ONNX Optimizer. OODTE adopts a straightforward yet powerful differential testing and evaluation methodology, which can be readily adapted for use with other compiler optimizers. Specifically, OODTE takes a collection of ONNX models, applies optimizations, and executes both the original and optimized versions across a user-defined input set, automatically capturing any issues encountered during optimization. When discrepancies in accuracy arise, OODTE iteratively isolates the responsible optimization pass by repeating the process at a finer granularity. We applied OODTE to 130 well-known models from the official ONNX Model Hub, spanning diverse tasks including classification, object detection, semantic segmentation, text summarization, question answering, and sentiment analysis. Our evaluation revealed that $9.2$% of the model instances either caused the optimizer to crash or led to the generation of invalid models using default optimization strategies. Additionally, $30$% of classification models and $16.6$% of object detection and segmentation models exhibited differing outputs across original and optimized versions, whereas models focused on text-related tasks were generally robust to optimization. In total, OODTE uncovered 15 issues—14 of which were previously unknown—related to optimizer crashes and accuracy regressions, impacting 9 of the 47 available optimization passes as well as the optimizer more broadly. All findings were reported to the ONNX Optimizer development team. OODTE demonstrates a simple but effective framework for validating AI model optimizers and with minimal work, is applicable beyond the ONNX ecosystem.
I Introduction
Deep Learning (DL), primarily implemented through Deep Neural Networks (DNNs), is applied across a wide range of domains, including autonomous vehicles and medical imaging. As these systems become more prevalent and are assigned increasingly complex tasks, the demand for their computational efficiency intensifies. To address this, newer and more powerful hardware devices and architectures are being developed, and effectively utilizing their performance depends on compilers and the optimizations they apply to deep learning models [13, 17]. Widely-used AI compilers and frameworks, such as Apache TVM [12] and MLIR [32], incorporate their own native sets of optimization passes for efficient use of hardware. These include standard techniques such as operator fusion and constant folding, as well as more aggressive approaches like reduced-precision arithmetic, including the use of fast-math optimizations, which may compromise numerical accuracy.
Models can also be represented using industry-standard formats, such as ONNX [8]. By applying transformations to models in these formats, it is possible to reduce model size and improve performance. For ONNX in particular, the ONNX Optimizer [4] serves as the default tool for performing graph-level optimizations. As part of the official ONNX toolkit, it focuses on simplifying model structure while aiming to preserve functional correctness, particularly in terms of accuracy. However, the degree to which these optimizations maintain model accuracy remains an open question.
We investigate the extent to which ONNX Optimizer transformations affect the functional correctness of DNN models. Ideally, such transformations should not introduce any deviations in model behavior. To this end, we propose the Differential Testing of the ONNX Optimizer (OODTE), a framework for performing differential analysis between original and optimized models. Using OODTE, we collected 130 models of various types, including classification, object detection, text comprehension and generation from the official ONNX Model Hub [7]. Each model was optimized using ONNX Optimizer and evaluated on standard benchmark datasets to compare accuracy before and after optimization. In cases where discrepancies were observed, we conducted pass-by-pass optimizations to isolate the specific transformations responsible for the faults.
We focus on the ONNX Optimizer due to its impact and popularity (approximately $700$ GitHub stars) as a model optimization tool. In addition, it is integrated into the main ONNX repository and serves as a foundation for other widely adopted utilities, such as the ONNX Simplifier [5], which has received over 4,000 GitHub stars. However, existing ONNX Optimizer tests primarily focus on toy models rather than real-world models, primarily used for unit testing purposes. We aim to address this gap by introducing a differential testing methodology through OODTE. Our approach is not limited to ONNX; it is generalizable to other AI compilers and optimization frameworks.
In total, we observed output differences between the original models and their and optimized counterparts in $34$ out of the $130$ models ($26.1$%). Additionally, we encountered issues with $12$ model instances ($9.2$%) while using the primary (fuse and eliminate) optimizer passes - with either the optimizer crashing or generating an invalid model. Through detailed analysis, we identified $15$ distinct bugs, of which only one had been previously reported—and even then, only partially and without a conclusive explanation. In that case, we were able to provide substantial additional information not previously documented. All identified bugs were reported to the official ONNX Optimizer repository, accompanied by detailed descriptions, so that the optimizer developers can further examine and resolve them.
Overall, we present the following contributions:
1. OODTE  - a differential testing tool to examine the effects of the ONNX Optimizer to DNNs.
2. A survey of 130 models optimized using the ONNX Optimizer, presenting the effects of the optimizations on model correctness in detail.
3. A set of faults detected using OODTE - reported to the official ONNX Optimizer repository.
II Background
II-A Deep Learning Models
Deep neural networks (DNNs) are widely used for prediction tasks across domains. We focus on three common categories:
II-A1 Image Classification
Image classification, one of the earliest DNN applications, typically uses Convolutional Neural Networks (CNNs) to extract image features via learned convolution filters. These models output a ranked list of top-K predictions, each paired with a confidence score, typically normalized using a Softmax function [46].
II-A2 Object Detection and Semantic Segmentation
Object detection extends classification by identifying object locations using bounding boxes. Semantic segmentation further refines this by labeling each pixel with a class (e.g., road, pedestrian), offering dense scene understanding. Both rely heavily on convolutional operations and are critical in areas like autonomous driving.
II-A3 Text Comprehension and Generation
These models handle tasks such as summarization, recognition, and sentiment analysis. Text generation models, especially transformers [66], use contextual representations to iteratively predict the next token using a sliding window approach. This generative process continues until a stopping condition is met.
II-B Graph Optimizations
For deep neural networks (DNNs) to perform effectively on complex tasks, they must be optimized to operate more efficiently while maintaining their performance. A common approach to achieving this is optimizing the model graph to improve computational efficiency. This may involve techniques such as node pruning, node fusion, and constant folding, among others. Similar to conventional software, these optimizations are applied to DNNs, with AI compilers [12, 32] incorporating such techniques. Additionally, tools like the ONNX Optimizer, which we examine in this work, are specifically designed for this purpose.
II-C The ONNX Optimizer
The ONNX Optimizer is a graph-level optimization tool designed to improve the performance and portability of ONNX models by applying a series of transformation passes. It consists of a set of $47$ passes grouped into three main types: fusions, eliminations, and graph rewrites, with the purpose of improving model efficiency. Fusion passes merge operations (e.g., fuse_bn_into_conv, which merges BatchNorm and Conv nodes) to reduce overhead. Elimination passes remove redundant nodes, such as eliminate_deadend (removing nodes with unused outputs) and eliminate_nop_transpose (removing transpose nodes that do nothing). Graph rewrite passes modify graph structure for compatibility, such as rewrite_input_dtype, which adjusts input data types. By default, the optimizer applies only fusion and elimination passes; however, it also supports applying any subset of its $47$ transformation passes, based on its user-defined configuration.
II-D Model Output Comparison Methods
Based on the model and output type, OODTE uses a set of well-known and established methods in order to compare the outputs of the optimized model versions with their original counterparts. In particular:
Classification models:
We compare the top-K label predictions between original and optimized models using Kendall’s Tau Rank Correlation [28], which evaluates the similarity between ranked label lists, with scores ranging from -1 (inverse correlation) to 1 (perfect correlation).
Object Detection and Semantic Segmentation models:
Each prediction includes (1) class probabilities, (2) bounding box coordinates, and (3) a confidence score. For class probabilities, we apply a softmax and use Kendall’s Tau to compare rank similarity. For bounding boxes and confidence scores, we use standard metrics: (I) IoU for bounding box overlap, (II) F1 Score for overall detection balance, (III) Average Precision (AP) for precision across thresholds, (IV) Average Recall (AR) for recall across thresholds, (V) mAP for aggregate detection and localization accuracy. These metrics are computed at IoU thresholds of 0.5, 0.75, and 0.9, with the original model as the reference.
Text generation models:
We convert token outputs to plain text and use the BLEU Score [45] to compare the optimized model’s output against the original, assessing n-gram overlap and overall similarity.
Binary sentiment analysis models:
We measure the percentage of prediction label differences between original and optimized models to evaluate the impact of optimization on classification consistency.
III System Architecture & Methodology
Figure 1: Architecture of OODTE  with the following primary modules: (1) Model Orchestrator, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module.
The architecture of OODTE with its different modules is shown in Figure 1. It has the following intuitive workflow: load models, optimize them, execute both the original and the optimized versions across the dataset and compare the results. If any discrepancies in model outputs are detected, OODTE  performs an iterative approach where each model under test is optimized with each ONNX Optimizer pass in isolation, and then the results are compared with the original model. This is done to localize the optimization pass responsible for causing an output difference. All results are stored in report files, both for reporting and localization purposes. In detail, OODTE  consists of the following primary modules: (1) Model Orchestrator, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module. OODTE  loads model configuration settings in two ways: (1) via local configuration file, and (2) by fetching configuration via the ONNX repository from which the models are loaded from. This is managed by the Model Orchestrator Module. The modules within OODTE  in Figure 1 is explained in the following Sections.
III-A Model Orchestrator Module
The Model Orchestrator Module is responsible for all tasks related to fetching, loading (including model configuration) and preparing models for optimization and execution. The module supports two loading options for DNN models: (1) local loading of ONNX models, and (2) loading of an online repository. At the moment, OODTE supports the official ONNX Model Hub [7], however, as it utilizes the official Hub API, it can trivially support third party ONNX model repositories. In addition, the module supports fetching and loading of multiple models simultaneously, while it also performs model caching when fetching from a model hub, as well as checksum verification. Once each model and its respective configuration are properly loaded, they are propagated to the Optimizer Module and the Runner Module.
III-B Optimizer Module
The Optimizer module is responsible for the generation of optimized models, using the ONNX Optimizer at its heart. It supports two settings: (1) full optimization, applying all the ONNX Optimizer passes to the model under test, and (2) partial optimization, where only a select set of passes is applied to the model. It is worth noting that in (1), by optimizer design, only the fuse and eliminate passes are applied as part of the default optimization bundle in the ONNX Optimizer, leaving out a number of secondary passes (e.g., ‘rename_input_output‘). OODTE supports the utilization of these passes for testing, in order to explore all optimizer operations. This is useful across OODTE pipeline, when faulty behavior is discovered across model evaluation and the models presenting the faults need to be examined in depth. OODTE isolates such cases after a full pass optimization exeuction and performs per-pass optimization generation for that purpose.
Once a model is optimized with either setting, it then gets propagated to the Runner model, in order to perform inference.
III-C Runner Module
The Runner model does precisely what its name suggests: executes the models under test. In particular, it receives both the original and the optimized model under test, as well as a dataset that will be used as input for both model versions. It then performs model inference, using ONNX Runtime [2], combined with the necessary model preprocessing, while utilizing the model configuration loaded in the Model Orchestrator component. OODTE supports different types of datasets, based on the type of models being examined - from images to IMDB reviews. It can load the inputs from local files, but also fetch the dataset from well-known packages, such as the datasets [3] package.
To enable the use of large-scale datasets and extract results from models that are larger and more computationally demanding—especially when operating on low-end devices—the Runner Module supports execution across a dataset in parts. In particular, a run can be divided into $N$ chunks, where each chunk consists of a fraction, $1/N$th, of the dataset. Any remaining data not evenly divisible is also included in the final chunk of the run. This process applies to both the original and the optimized model versions, with the results extracted and propagated to the Metrics Comparison Module for subsequent evaluation and reporting. By following this approach, we were able to conduct our survey on two consumer-grade devices without relying on a compute cluster to obtain results, as described in Section V.
III-D Metrics Comparison Module
Once the run completes per-chunk for both the original and the optimized model, the results are propagated to the Metrics Comparison Module. This module performs comparison against the results, by selecting the appropriate comparator based on the model type. For classification types, OODTE supports first-label comparison, but also calculates the Kendall Rank Correlation Coefficient across the $top-K$ predictions - commonly known as Kendall’s Tau [28]. For Object Detection models, it utilizes the same methods of classification for labels and ranks, but also applies Intersection over Union (IoU [67]), Precision, Recall, and F1 Score [15] given an IoU threshold. For text generation models (e.g., GPT-2), it utilizes BLEU [45] score in order to compare the text generated out both the original and the optimized model, as well as direct tensor comparison for cases of question answering (Q&A). In addition OODTE is build in a manner that new comparators can be added, in cases that a new category of models is tested.
Once a method is selected, the Metrics Comparison Module calculates the metrics based on model type and generates a report with all results, in JSON format. As mentioned earlier, if a difference is detected across the original and the optimized model, then OODTE retrofits the reported data to the Optimizer Module. Then, the Optimizer Module generates model optimizations per-pass, reporting how each pass behaves.
IV Related Work
Conventional compiler testing is a domain that has been explored [10, 73, 16], involving differential testing [48, 65], as well as including compiler optimizations [11, 27, 19]. However, in terms of testing AI tools and compilers, most of the existing research has focused on (1) the direction of compiler fuzzing and (2) examining context-specific AI tools and APIs, such as DL framework converters.
In terms of compiler fuzzing, HirGen [43] is a utility for Deep Learning compiler fuzzing, focusing on the high-level model intermediate representation (IR) of TVM. TVMFuzz [1] allows generic-purpose TIR fuzzing on Apache TVM, detecting primarily crash and bit manipulation bugs. Tzer [38] is a coverage-guided fuzzer that applies modifications on models compiled in Apache TVM utilizing low-level IR (Tensor IR). GenCoG [69] focuses on generating valid and expressive TVM computation graphs by defining a domain specific language for constraint definition, as well as concolic constraint solving. NNSmith [37] focuses on valid DNN graph generation, usable to examine the compiler for bugs similarly with traditional grammar-based fuzzing techniques (e.g., [73, 16]), and MLIRSmith [68] integrates NNSmith’s fuzzing capabilities in MLIR [32] randomly. Suo et al. [60] also follow an operation dependency approach to guide the fuzzing process in MLIR. DeepDiffer [35] applies low-level IR fuzzing in a coverage-guided manner and comparing results across different DL compilers. SYNTHFUZZ [34] combines grammar-based fuzzing with custom mutation synthesis in order to effectively fuzz MLIR dialects. Li et al. [33] propose a methodology to uncover hardware-related faults by generating fault-sensitive, adversarial tests. Finally, Su et al. [59] fuzz DL compilers in the context of dynamic features, detecting bugs in PyTorch and its underlying tensor compiler.
Regarding testing techniques for context-specific AI tools and APIs, CRADLE [47] performs an execution analysis on the model graph in order to detect faults, while LEMON [70] utilizes the metrics used by CRADLE and applies mutation testing in order to identify issues in the model. FAME [57] is based on LEMON but enables API mutation and optimizations for layer and weight mutation, guided by the detection of inconcistency faults. Similarly, Audee [20] aims to detect logical, not-a-number bugs, and crashes by applying an exploratory approach in combination with mutation testing. Finally, Louloudakis et al. propose FetaFix [42] - a utility that allows automatic fault detection in model converters, involving ONNX as an intermediate model representation.
Overall, limited work has been done in the direction of testing AI compiler optimizations. Choi et al. [14] have conducted an empirical study towards optimizer comparisons, however they do not consider graph-based optimizations in the post-training phase such as the ONNX Optimizer) as they focus on stochastic optimizers focusing primarily on the model training process, such as Adam [29]. DiffChaser [72] explores the application of differential testing across optimized and unoptimized DNN model versions. Louloudakis et al. propose DeltaNN [41], a utility to explore model robustness across different configurations - including but not limited to compiler optimizations. Xiao et al. [71] propose a metamorphic testing methodology for AI compiler bug detection, including issues related to optimizations.
To the best of our knowledge, OODTE is the first that thoroughly tests the ONNX Optimizer (and AI optimizers in general), using differential testing in order to detect faults. The closest existing work to our methodology, is DiffChaser [72], DeltaNN [41] and FetaFix [42]. However, these tools considerably differentiate from our contribution. In particular, DiffChaser attempts to detect problems related to AI optimization techniques (such as quantization) related to the models under test, not exploring potential faults introduced from bugs in the optimizer. DeltaNN applies differential testing, but focuses on uncovering behavioral inconsistencies in the models under test when deployed using different configurations, rather than focusing on compiler issues. FetaFix, on the other hand, aims to uncover issues related to converters using differential testing and involving ONNX. Its approach however does not consider model optimizations or AI optimizer tools.
V Experiment Setup
V-1 The Experiment Process
We evaluated all ONNX Optimizer passes (as detailed in Section II-C). For each original model under test, we first applied the default setting of the ONNX Optimizer, which applies all fusion and elimination passes at once, generating the optimized model. We then compared the outputs of the original and optimized models according to the model type, as described in Section III-D. For optimized models presenting output differences, we performed optimization using only a single optimization pass, resulting in $47$ such single-pass optimized models. To identify which pass causes the problematic behavior, we compared the outputs of each single-pass model against those of the original model. It is worth noting that OODTE is compatible with new or custom passes as well. Furthermore, our methodology generalizes to other optimization frameworks (e.g., Apache TVM), provided they support both bundled and individual pass applications.
V-2 DNN Models
For our experiment set, we sourced $130$ model instances from the ONNX Model Hub [7]. In particular, we selected $93$ models utilized for classification tasks (e.g., MobileNet [24], EfficientNet-Lite4 [63] DenseNet121 [25], and InceptionV2 [61]), $30$ models used for object detection and semantic segmentation tasks (e.g., RetinaNet [36], Mask/Faster R-CNN R-50-FPN [21, 54], and SSD [39]), and $7$ transformer models [66] utilized for text comprehension and generation tasks, such as question answering (Q&A), (BERT-SQuAD [18]), sentiment analysis (RoBERTa [40]), text summarization (T5 [50]) and generic-purpose text generation (GPT-2 [49]), with $4$ models focusing on Q&A, while one model is utilized for each of the other text comprehension and generation tasks. We utilized models of a variety of opsets based on hub availability, ranging from $7$ (the lowest compatible with the latest ONNX Runtime version by the time of this publication - v1.21.11.21.11.21.11.21.1) up to $12$ (the highest provided from ONNX Model Hub via their API). Our methodology, however, is applicable to any ONNX model type and is not constrained by opset versions or tied to a specific model repository.
V-3 Computational Resources
We ran our models on mid-range systems with Intel i5 and i7 CPUs, reflecting the typical hardware used by developers during model development and fine-tuning. Our method is generalizable to larger models, limited only by ONNX model size constraints, and can be extended to high-end systems with a trivial extension for hardware acceleration (using the ONNX Runtime API [2], or an AI compiler such as Apache TVM [12]).
V-4 Datasets
We selected datasets appropriate to each model task. For classification and object detection, we used the ILSVRC2017 [55] validation dataset, containing $5500$ images, commonly used for benchmarking image recognition models. For text comprehension and generation, we used SQuAD-v2 [51], which includes $11873$ questions and contexts, to evaluate Q&A tasks and general-purpose transformer behavior. Additionally, for summarization and sentiment analysis, we used the IMDB dataset [44, 64], containing $25000$ entries.
To accurately detect differences across model types, we employed comparison methods appropriate for each task. Specifically, we used the following approaches:
  1. 1.
    For image recognition models performing classification tasks, we utilized Kendall Rank Correlation Coefficient [28] (KT), measuring differences for top-K prediction labels of each dataset input image, for $K=1,5$, and $10$. If the KT value was found to be $<1$ for each input, we counted this input as different.
  2. 2.
    For image recognition models performing object detection and semantic segmentation, we calculated precision, recall, and F1 score for each dataset input, and extracted their average, combined with different Intersection-over-Union (IoU) thresholds ($IoU=0.5,0.75$, and $0.9$), as well as IoU, mean Average Precision (mAP), along with average precision and recall. We considered the original model output as the base, and the optimizer model the one under test for our metric calculations, and the higher the values observed, the better the optimized model aligns with the original.
  3. 3.
    For text generation models performing question answering and summarization, we utilized BLEU score [45] across output texts for each input, monitoring instances where $BLEU\_Score<1$ for each input.
  4. 4.
    For sentiment analysis models, we compared each output value consisting of ${0,1}$ across the original and the optimized versions, registering any differences observed.
VI Results
Following the execution of the full set of experiments11footnotemark: 1, OODTE successfully identified $15$ bugs, reported across $11$ separate GitHub issues [9]. A complete list of these issues, along with the associated models, is shown in Table I. Of the 15 bugs identified, $14$ were entirely new (not previously reported on the ONNX Optimizer issue tracker), while one was a previously reported issue that lacked sufficient detail, including model information and reproducibility steps. We provided comprehensive information and full reproducibility details for this issue.
Within the ONNX optimizer, the bugs identified were primarily associated with $9$ of the $47$ passes. However, not all passes are equally important. By default, the optimizer prioritizes passes related to fuse and eliminate operations. As a result, $8$ of the $11$ reported GitHub issues are considered more critical since they belong to the optimizer’s core (default) strategy. Among these $8$ issues, $6$ are linked to specific passes (e.g., fuse_bn_to_conv, which fuses batch normalization nodes with convolutions), while the remaining 2 issues were more general and related to the optimizer’s overall usage. We analyze and present the impact of these faults both on the optimization process and on the correctness of the optimized models in the following Sections.
TABLE I: Overview of the issues identified using OODTE, detailing the responsible optimization pass, the detected error, its type (crash or accuracy), and the models in which the error was observed, along with their associated opsets (op.).
# Pass Error Type Models
1 lift_lexical_references Removes used value references. Crash EfficientNet-Lite4 (op. 11), SSD (op. 12)
2 split_init List index out of range. Crash YOLOv3 (op. 10, 12)
3 split_init split_predict Field is required but missing. Crash DenseNet121 (op. 7-9), YOLOv2 (op. 10), SSD (op. 12), Tiny-YOLOv3 (op. 11) YOLOv3 (op. 10, 12)
4 split_predict Required inputs are missing. Crash SSD (op. 12), DenseNet121 (op. 7), InceptionV2 (op. 6-9)
5 split_predict Bad spec for node. Crash YOLOv3 (op. 10), DenseNet121 (op. 7) InceptionV2 (op. 6-9)
6 split_predict Input element number mismatch. Crash FCN ResNet-50-int8 (op. 12)
7 eliminate_shape_gather Indices value not equal to node dims size Crash Faster R-CNN: R-50-FPN (op. 10)
8 All Passes Implicit ONNX format version change. Accuracy Models of cases $1$-$11$ in Table II
9 rewrite_input_dtype Output name used multiple times. Crash GPT-2 (op. 10)
10 eliminate_ consecutive_ idempotent_ops Incompatible shape error. Crash (Run) Mask R-CNN/Faster R-CNN: R-50-FPN-int8 (12), R-50-FPN-qdq (op. 12), R-50-FPN-fp32 (op. 12)
11 eliminate_ duplicate_ initializer Erase output assertion failed. Crash Faster R-CNN: R-50-FPN-qdq (12)
12 fuse_bn_into_conv Duplicate initializer introduced. Crash (Run) RetinaNet (op. 9)
13 eliminate_shape_gather eliminate_nop_reshape Added initializers not used by the model. Warning Faster R-CNN: R-50-FPN-int8 (op. 12)
14 All Passes Invalid input scale. Crash VGG-16-BN (op. 7), VGG-19-BN (op. 7), ResNet101_DUC_HDC (op. 7)
15 fuse_bn_into_conv Bad graph structure. Accuracy Efficient-Lite4 (op. 11), SSD (op. 12)
VI-A Optimizer Crashes
Out of the $15$ detected bugs, $12$ bugs were related to model crashes, $2$ bugs resulted in output differences between the original and the optimized model, and $1$ bug introduced redundant initializers in the model that produced runtime warnings. In particular, we observed crashes were caused by $8$ out of the $47$ passes ($17$%), and occurred for specific models, with object detection models being more prone to errors from these $8$ passes. In particular, R-CNN models, such as Faster R-CNN R-50-FPN [54], RetinaNet [36], VGG-16, VGG-19 [58], as well as YOLOv3 [53], presented crashes by using the optimizer, in a range of opset instances and in passes related to the modification of gather nodes, and the elimination of idempotent operations. In total, we observed $31$ model instances where either the optimizer crashed upon processing them, or the optimized model was problematic, leading to a runtime crash. Some secondary passes are known to be unstable (e.g., split_init and split_predict describe in their documentation that they not strictly comply with the ONNX format), and therefore we do not focus on them as sources of faults in the optimization process. Overall, $12$ out of the $31$ model instances ($9.2$%) that were associated with a crash occurred while using passes related to fuse and eliminate operations, which are part of the default bundle of optimizations in the ONNX Optimizer and are therefore of particular significance.
As shown in Table I, crash faults can be attributed to three main causes: (1) model input issues (cases #$4$ and #$6$), (2) incorrect use of initializers (cases #$12$ and #$13$), and (3) issues associated with node properties such as duplicate names, obsolete fields, and incorrect dimensions (all other cases). Overall, object detection models were more prone to optimizer crashes compared to classification models. Due to their increased structural complexity and larger size, they are more likely to exhibit errors related to node properties and initializers, which account for the vast majority of crash faults. In contrast, only one issue was observed in the most complex model category — text generation models, that was related to model input in GPT-2. However, upon further inspection, we found that the transformation passes of the ONNX Optimizer had limited impact on text generation models — the majority of optimization passes left the models unchanged, and those that did make modifications affected a small proportion of nodes with negligible effect on outputs.
VI-B Output Label Differences in Optimized Classification & Object Detection Models
Out of the $130$ optimized models tested, we observed unexpected output differences between the original and optimized models. These differences manifested as (1) label differences for classification models (e.g., ResNet), (2) label and/or bounding box differences for object detection models (e.g., YOLOv3), and (3) textual output differences in token sequences or predicted spans for text generation tasks. Since the ONNX Optimizer transformations are graph-based and do not alter numeric precision, output differences are not expected. Among the $93$ classification models tested, $28$ ($30.1$%) exhibited discrepancies in output labels. Similarly, $5$ out of $30$ ($16.6$%) object detection models presented output differences. The text generation models, on the other hand, were robust with only one instance (GPT-2) exhibiting negligible differences ($<0.001\%$) while all others showing identical results across the original and the optimized model versions. The differences primarily affected the top-5 and top-10 predictions, with no changes in the top-1 predictions, except for the SSD model, which showed a $2.9\%$ difference in top-1 prediction. Although top-1 accuracy is often prioritized, it is important to recognize that changes in lower-ranked predictions can also indicate issues. Our methodology identified these inconsistencies, which have been reported to the ONNX Optimizer developers.
Figure 2: Illustrative example of problematic label classification and object detection between original unoptimized (blue) and optimized (orange) models. The bounding box location is affected for both bird instances, while the object label prediction is affected by model optimization on just the right bird instance, changing the label from bird to jet.
TABLE II: Output label accuracy differences across classification (CL) and object detection (OD) models, showing the type, opset versions, and accuracy differences for top-1, top-5, and top-10 output predictions, between the original and optimized (using the default ONNX Optimizer setting) versions of each model, along with the issue related to the observed differences.
# Model Type Opset(s) Top 1 Top 5 Top 10 Issue
1 AlexNet [30] CL 7-9 0% 0% 0.02% Format Version Change
2 DenseNet121 [25] CL 7-9 0% 0% 0.02% Format Version Change
3 GoogleNet [62] CL 7-9 0% 0% 0.04% Format Version Change
4 Inception-1 [61] CL 7-9 0% 0% 0.25% Format Version Change
5 Inception-2 [61] CL 7-9 0% 0.02% 0.05% Format Version Change
6 MobileNet v2 [24, 56] CL 7-9 0% 0% 0.02% Format Version Change
7 ResNet18 [22] CL 7 0% 0.02% 0.02% Format Version Change
8 ResNet34-v2 [23] CL 7 0% 0.02% 0.02% Format Version Change
9 ResNet50-caffe2 [22] CL 7-9 0% 2% 13% Format Version Change
10 ShuffleNet-v1 [74] CL 7-9 0% 0.7% 1% Format Version Change
11 SqueezeNet 1.0 [26] CL 7-9 0% 0% 0.02% Format Version Change
12 EfficientNet-Lite4 [63] CL 11 0% 0.04% 1% Graph Malformation
13 SSD [39] OD 12 2.9% 4.3% 6% Graph Malformation
Overall, we detected two primary issues related to output label differences between the original and optimized models:
Implicit ONNX format version upgrade:
The ONNX Optimizer silently upgrades models from ONNX format version v3 to v4, without any warning or indication. This behavior impacts models using opset versions 7, 8, and 9— that are older opsets but still supported in the current ONNX Runtime (v1.21.1). Despite being outdated, these opset versions are still prevalent, representing approximately 41% of the models in the ONNX Model Hub. While not all models within this opset range are affected, we observed a significant number that exhibited differences between their original and optimized versions. We have reported this issue to the ONNX Optimizer developers, who have acknowledged it and are investigating further.
Problematic fusion pass:
The pass fuse_bn_to_conv fuses Batch Normalization and Convolution operators within the model. In particular, the pass attempts to fuse BatchNormalization nodes (which do not update their mean or variance during inference) with their preceding Conv nodes by adjusting the weights and biases of the Conv node according to the properties of the BatchNormalization node. We observed that this pass was problematic in a number of cases, resulting in both a crashing bug and an issue where redundant nodes are introduced into the model. In the latter scenario, instead of properly merging the nodes resulting in a Conv node with adjusted weights and biases, the pass generates multiple nodes that represent the Batch Normalization formula. This behavior is illustrated on the right side of Figure 3. This formula representation is problematic as several models produced different output labels after optimization using this pass compared to their original versions. Although another user previously reported a similar issue, no details on the reproduction steps or affected models were provided. We submitted a reproducible issue to the ONNX Optimizer developers, including all relevant information and behavior details.
Table II shows label differences in top-K outputs for classification and object detection models, highlighting the causes. As an example, Figure 2 illustrates how an optimized model (orange) can misclassify and alter bounding boxes compared to the original model (blue).
Figure 3: Incorrect model graph transformation caused by the fuse_bn_to_conv pass in EfficientNet-Lite-4. The nodes affected are highlighted. The result should produce a Conv node with updated weights and biases, but the pass generates multiple nodes that represent the BatchNormalization formula.
TABLE III: Object detection metric differences across the original and optimized versions of object detection models, using the default ONNX Optimizer setting. The table presets the name, opset, average F1 score, precision, recall, mean Average Precision (MAP) and average Intersection over Union (IoU scores), for different IoU thresholds ($50$%, $75$%, and $90$%).
Model Op. Avg. F1 Avg. Prec. Avg. Recall mAP Avg. IoU
SSD [39] 12 95.28% 95.28% 95.28% 86.62% 97.37%
Tiny YOLOv3 [53] 11 99.56% 99.56% 99.56% 90.57% 99.61%
YOLOv2 [52] 9 99.95% 99.95% 99.95% 90.85% 88.50%
YOLOv3 [53] 10 98.83% 98.83% 98.83% 86.66% 98.78%
YOLOv3 [53] 12 98.83% 98.83% 98.83% 86.65% 98.78%
(a) IoU threshold of 50%.
Model Op. Avg. F1 Avg. Prec. Avg. Recall mAP Avg. IoU
SSD [39] 12 95.27% 95.27% 95.27% 86.6% 97.37%
Tiny YOLOv3 [53] 11 99.36% 99.36% 99.36% 89.97% 99.61%
YOLOv2 [52] 9 99.97% 99.97% 99.97% 90.87% 88.60%
YOLOv3 [53] 10 97.87% 97.87% 97.87% 85.71% 98.78%
YOLOv3 [53] 12 97.87% 97.87% 97.87% 85.71% 98.78%
(b) IoU threshold of 75%.
Model Op. Avg. F1 Avg. Prec. Avg. Recall mAP Avg. IoU
SSD [39] 12 95.27% 95.27% 95.27% 86.6% 97.37%
Tiny YOLOv3 [53] 11 99.34% 99.34% 99.34% 89.97% 99.61%
YOLOv2 [52] 9 99.97% 99.97% 99.97% 90.87% 88.60%
YOLOv3 [53] 10 97.8% 97.8% 97.8% 85.69% 98.78%
YOLOv3 [53] 12 97.8% 97.8% 97.8% 85.69% 98.78%
(c) IoU threshold of 90%.
VI-C Bounding Box Differences in Object Detection Models
We observed notable differences in bounding box outputs between the original and optimized versions of object detection models. The results are presented in detail in Table III with the following main findings:
YOLOv3 Models:
Both YOLOv3 models (op.set $10$ and $12$) achieve high average IoU, F1 score, precision, and recall—all above $98$%. Although the optimized model occasionally misplaces bounding boxes, these errors do not cause misclassifications, as none were found in the top-10 label predictions.
YOLOv2 Model:
YOLOv2 shows lower IoU values, likely because it uses multiple bounding boxes per detected object. Small differences between the original and optimized models accumulate, lowering IoU. However, high F1, precision, and recall indicate that detection performance remains largely unaffected, as demonstrated by the bounding boxes corresponding to the bird on the left in Figure 2.
SSD Model:
In contrast, SSD exhibits a $4.8$% drop in F1, precision, and recall between original and optimized models, resulting in 2.92.92.92.96666% misclassification rates despite maintaining a high IoU ($>97$%). This behavior is illustrated in Figure 2 by the bounding boxes associated with the bird on the right.
All models show a substantial mAP difference of up to $14.31$% between original and optimized model versions. This suggests the optimizer may cause certain classes to be missed, affecting both object detection and semantic segmentation. Consequently, even if detected objects are classified correctly, some objects in the scene might be overlooked, which could significantly impact real-world applications.
VI-D Optimizer Warnings
We observed two issues related to the addition of redundant elements, specifically unused initializer entries introduced in the optimized models, associated with the eliminate_shape_gather and eliminate_nop_reshape passes. When executed with ONNX Runtime, these entries were ignored, triggering a warning in the format: Removing initializer ‘INITIALIZER_ID’. It is not used by any node and should be removed from the model. Upon manual inspection, we found no significant differences in model size. However, this issue warrants further investigation for potential side effects. We have reported it to the optimizer developers for additional review.
VII Tool Usability
OODTE  can test whether the ONNX Optimizer generates optimized models without compromising output label correctness. Currently, OODTE supports differential testing for classification, object detection, and text generation models. However, OODTE can also be easily extended to support new model types, as long as (1) the models can be compiled into the ONNX format for optimization and (2) a comparator is provided to compare the results between the base and optimized models for the new model type. In addition, OODTE can fetch models and datasets automatically, making it adaptable for use in a Continuous Integration pipeline. This allows OODTE to test the ONNX Optimizer whenever new functionality (e.g., new passes) is added. By default, OODTE loads models from the official ONNX Model Hub [7], which was used in our experiments. However, by configuring OODTE to use a custom ONNX Hub that follows the ONNX API specification for model repositories [6], it can fetch and use models without further modification. The same flexibility applies to datasets. OODTE can fetch datasets using packages like the datasets [3] package or from local storage.
Finally, while OODTE applies a differential testing methodology focused on the ONNX Optimizer, this approach is generalizable and can be easily adapted to other AI optimizers or compilers that meet two primary requirements: (1) they provide a runtime for model execution, and (2) they expose an API that includes all the passes applied during model optimization, both as a bundle and as independent passes.
VIII Threats To Validity
We list the external and internal threats to validity in our evaluation below.
VIII-A External Threats
There are three primary external threats to the validity of our findings.
(1) Model Selection: Our evaluation included 130 models sourced from the official ONNX Model Hub, covering widely used categories such as classification, object detection, and text comprehension and generation. These models were selected for their maturity, accessibility, and broad adoption across academic and industrial domains. While we aimed to create a representative and diverse experimental set, the selection is not exhaustive. Models with opset versions below 7 were excluded due to incompatibility with the current ONNX and ONNX Runtime versions, which may limit our findings to more recent models. Additionally, to keep the evaluation tractable, we did not include certain model types—such as diffusion models and reinforcement learning—which may behave differently under optimization. Although our tool supports any ONNX-compliant model hub, the exclusion of these model types presents a potential threat to the generalizability of our results. We plan to address this limitation in future work.
To keep the experiments tractable, we did not evaluate model types such as diffusion models and reinforcement learning in this work; however, we plan to include them in future studies.
(2) Dataset Selection: Our experimental evaluation relies on widely used, real-world benchmark datasets tailored to the specific tasks of the models under test. While this approach ensures relevance and practical applicability—e.g., using ILSVRC [55] for image classification and object detection models like ResNet [22] and GoogleNet [62], and leveraging established HuggingFace datasets [3] for NLP tasks—it may limit generalizability. The use of popular datasets could introduce bias toward well-studied scenarios and may not reflect performance on less common or emerging tasks.
(3) Devices Used: Our experiments were conducted on two hardware platforms: an Intel i5 and an Intel i7 processor. While we verified consistency by (a) randomly sampling models of each type and running them on both devices with consistent results, and (b) referencing prior work by Louloudakis et al. [41], which found that optimization does not affect accuracy across processing devices, this still presents a potential threat to validity. The limited range of hardware may not capture performance or accuracy variations that could arise on other architectures, such as GPUs or edge devices, and may limit the generalizability of our findings.
VIII-B Internal Threats
We identify the following internal threat to validity in our findings.
Optimization Pass Ordering: The order in which optimization passes are applied can significantly affect model performance and behavior, as highlighted in prior work [31]. This is a well-known challenge in compiler research. While this work does not explore the impact of pass ordering in depth, we adopted two evaluation modes in OODTE: (1) the default ONNX Optimizer configuration, which includes commonly used passes such as fuse and eliminate, and (2) a configuration where all 47 available optimization passes were applied atomically. Although this strategy does not account for potential issues arising from specific pass combinations or orderings, it has proven effective in uncovering correctness issues in practice. Nevertheless, the omission of a systematic study on optimization ordering may limit the completeness of our evaluation.
IX Discussion
We summarize below our findings for the different model types in our experiments.
(1) Object detection models crash more frequently when optimized.
Most crashes during optimization occurred in object detection and semantic segmentation models, rather than in classification models. This is likely due to their greater architectural complexity—being larger and structurally more complex—which leads the ONNX Optimizer to apply more extensive transformations, increasing the likelihood of encountering corner cases that trigger crashes related to the ONNX Optimizer passes.
While text generation and comprehension models are also typically large, manual inspection revealed that the ONNX Optimizer applies only a limited set of optimizations to them. Most passes are skipped, as many of the layer patterns targeted by the optimizer (e.g., convolutions)—common in classification and object detection models—are not typically present in transformer-based architectures used for language tasks.
(2) Optimization can reduce the accuracy of object detection models.
While some object detection models remained relatively stable after optimization, others showed notable performance degradation. For example, the YOLOv3 model maintained high detection accuracy with no misclassifications, although it exhibited minor shifts in bounding box positions. YOLOv2 also showed stable performance, with slightly lower Intersection over Union (IoU) due to its use of multiple bounding boxes per object, but still achieved strong detection metrics overall.
In contrast, the SSD model experienced a significant decline in F1 score, precision, and recall, leading to higher misclassification rates despite retaining a high IoU. Across all models, we observed substantial drops in mean Average Precision (mAP), with reductions of up to 14.31%, suggesting that optimization may cause certain object classes to be missed. This raises concerns about the reliability of optimized models in object detection and segmentation tasks.
(3) Optimizations can affect label accuracy in classification models. We identified two primary causes for differences in label accuracy following optimization. (a) The first relates to the implicit and improper upgrade of the ONNX format version in models using older opset versions. A significant number of classification models were built on these legacy opsets, making them particularly susceptible to this issue. (b) The second involves the generation of problematic graph nodes, particularly in relation to the fuse_bn_into_conv pass. This pass, which fuses batch normalization into convolutional layers, has a strong impact on classification models and can lead to unintended changes in model behavior after optimization.
(4) Outputs from Text comprehension and generation models were not affected by optimization. From the $8$ model instances under test related to text comprehension and generation, we only found one crash issue related to GPT-2, as well as a very small number of samples (8) producing a BLEU score that indicates mismatch between the original and the optimized model, for the top-3 token selections of the model.
Overall, we find ONNX optimizations can result in output differences between the optimized and unoptimized model, with some model types being more susceptible to optimization errors than others. OODTE  was able to successfully isolate the optimization passes responsible for the output differences for different model types.
X Conclusion
We propose OODTE, a tool for the differential testing of the ONNX Optimizer. Using OODTE, we evaluated $130$ models across classification, object detection, and text generation tasks. We detected $9.2$% of model instances that crashed when using the primary optimizer strategies, while $30$% of classification models and $16.6$% of object detection models showed differences between the source and optimized versions, while text generation models remained robust. Additionally, OODTE identified $15$ previously unreported bugs related to crashes and accuracy degradation across $9$ optimizer passes, which we reported to the ONNX Optimizer developers.
XI Availability
For more information about OODTE, please contact the authors. We intend to release the tool’s source code in the near future.