### v3

# OODTE: A Differential Testing Engine for the ONNX Optimizer

Abstract.
With over 760 stars on GitHub and being part of the official ONNX repository, the ONNX Optimizer is the default tool for applying graph-based optimizations to ONNX models. Despite its widespread use, its ability to maintain model accuracy during optimization has not been thoroughly investigated. In this work, we present OODTE, a utility designed to automatically and comprehensively evaluate the correctness of the ONNX Optimizer. OODTE adopts a straightforward yet powerful differential testing, fault localization, and evaluation methodology, which can be readily adapted for use with other compiler optimizers. Specifically, OODTE takes a collection of ONNX models, applies optimizations, and executes both the original and optimized versions across a user-defined input set, automatically capturing any issues encountered during optimization. When discrepancies in accuracy arise, OODTE iteratively isolates the responsible optimization pass by repeating the process at a finer granularity. We applied OODTE to 130 well-known models from the official ONNX Model Hub, spanning diverse tasks including classification, object detection, semantic segmentation, text summarization, question answering, and sentiment analysis. Our evaluation revealed that 9.29.2% of the model instances either caused the optimizer to crash or led to the generation of invalid models using default optimization strategies. Additionally, 3030% of classification models and 16.616.6% of object detection and segmentation models exhibited differing outputs across original and optimized versions, whereas models focused on text-related tasks were generally robust to optimization. In total, OODTE uncovered 15 issues, 14 of which were previously unknown—related to optimizer crashes and accuracy regressions, impacting 9 of the 47 available optimization passes as well as the optimizer more broadly. All findings were reported to the ONNX Optimizer development team. OODTE demonstrates a simple but effective framework for validating AI model optimizers and with minimal work, is applicable beyond the ONNX ecosystem.
1. Introduction
Deep Learning (DL), primarily implemented through Deep Neural Networks (DNNs), is applied across a wide range of domains, including autonomous vehicles and medical imaging. As these systems become more prevalent and are assigned increasingly complex tasks, the demand for their computational efficiency intensifies. To address this, newer and more powerful hardware devices and architectures are being developed, and effectively utilizing their performance depends on compilers and the optimizations they apply to deep learning models (speedisallyouneed, ; CompilerGym, ). Widely-used AI compilers and frameworks, such as Apache TVM (tvm, ) and MLIR (lattner2020mlircompilerinfrastructureend, ), incorporate their own native sets of optimization passes for efficient use of hardware. These include standard techniques such as operator fusion and constant folding, as well as more aggressive approaches like reduced-precision arithmetic, including the use of fast-math optimizations, which may compromise numerical accuracy.
Models can also be represented using industry-standard formats, such as ONNX (onnxsite, ). By applying transformations to models in these formats, it is possible to reduce model size and improve performance. For ONNX in particular, the ONNX Optimizer (onnxoptimizer, ) serves as the default tool for performing graph-level optimizations. As part of the official ONNX toolkit, it focuses on simplifying model structure while aiming to preserve functional correctness, particularly in terms of accuracy. However, the degree to which these optimizations maintain model accuracy remains an open question.
We investigate the extent to which ONNX Optimizer transformations affect the functional correctness of DNN models. Ideally, such transformations should not introduce any deviations in model behavior. To this end, we propose the Differential Testing of the ONNX Optimizer (OODTE), a suite that detects and localizes faults in the ONNX Optimizer.
Problem Being Solved
OODTE solves the problem of detection and localization of faults associated with the ONNX Optimizer, primarily in the direction of output label inconsistencies but also compiler crashes and miscompilations. In particular, OODTE employs a per-pass fault localization strategy combined with differential testing between original and optimized models of various types to identify and localize issues arising from specific passes in the ONNX Optimizer.
Case Studies
To evaluate the effectiveness of OODTE, we collected 130 models as case studies, performing different tasks, such as classification, object detection, text comprehension, and text generation. All models were obtained from the official ONNX Model Hub (onnxmodelhub, ). Each model was optimized using the ONNX Optimizer and assessed on standard benchmark datasets to compare accuracy before and after optimization. When discrepancies were detected, we applied pass-by-pass optimization to isolate the specific transformations responsible for the faults.
We focus on the ONNX Optimizer due to its impact and popularity (over 760 GitHub stars) as a model optimization tool. In addition, it is integrated into the main ONNX repository and serves as a foundation for other widely adopted utilities, such as the ONNX Simplifier (onnxsimplifier, ), which has received over 4,000 GitHub stars. However, existing ONNX Optimizer tests primarily focus on toy models rather than real-world models, primarily used for unit testing purposes. With OODTE, we aim to enhance the effectiveness of fault localization and detection in the optimizer. However approach is not limited to ONNX; it is generalizable to other AI compilers and optimization frameworks.
By conducting experiments on our case studies, we observed output discrepancies between the original and optimized models in $34$ of $130$ cases (26.126.1%). We also encountered issues with $12$ model instances (9.29.2%) using the primary optimizer passes (fuse and eliminate) where the optimizer crashed or produced invalid models. Through analysis, we identified 15 distinct bugs, only one of which had been partially reported before (without a full explanation or reproduction steps). For this case, we provided additional insights. All bugs were reported to the official ONNX Optimizer repository with detailed descriptions for further resolution. OODTE effectively detected faults due to incorrect value references, input/output mismatches, and graph-structure inconsistencies introduced by the ONNX optimizer. These manifested as (1) optimizer crashes, (2) malformed models, and (3) valid models with output label inconsistencies compared to their unoptimized counterparts.
2. Background
2.1. Deep Learning Models
Deep Neural Networks (DNNs) are widely used for prediction tasks across domains. We focus on three common categories:
2.1.1. Image Classification
Image classification, one of the earliest DNN applications, typically uses Convolutional Neural Networks (CNNs) to extract image features via learned convolution filters. These models output a ranked list of top-K predictions, each paired with a confidence score, typically normalized using a Softmax function (softmax, ).
2.1.2. Object Detection and Semantic Segmentation
Object detection extends classification by identifying object locations using bounding boxes. Semantic segmentation further refines this by labeling each pixel with a class (e.g., road, pedestrian), offering dense scene understanding. Both rely heavily on convolutional operations and are critical in areas like autonomous driving.
2.1.3. Text Comprehension and Generation
These models handle tasks such as summarization, recognition, and sentiment analysis. Text generation models, especially transformers (attention, ), use contextual representations to iteratively predict the next token using a sliding window approach. This generative process continues until a stopping condition is met.
2.2. Graph Optimizations
For deep neural networks (DNNs) to perform effectively on complex tasks, they must be optimized to operate more efficiently while maintaining their performance. A common approach to achieving this is optimizing the model graph to improve computational efficiency. This may involve techniques such as node pruning, node fusion, and constant folding, among others. Similar to conventional software, these optimizations are applied to DNNs, with AI compilers (tvm, ; lattner2020mlircompilerinfrastructureend, ) incorporating such techniques. Additionally, tools like the ONNX Optimizer, which we examine in this work, are specifically designed for this purpose.
2.3. The ONNX Optimizer
The ONNX Optimizer is a graph-level optimization tool designed to improve the performance and portability of ONNX models by applying a series of transformation passes. It consists of a set of $47$ passes grouped into three main types: fusions, eliminations, and graph rewrites, with the purpose of improving model efficiency. Fusion passes merge operations (e.g., fuse_bn_into_conv, which merges BatchNorm and Conv nodes) to reduce overhead. Elimination passes remove redundant nodes, such as eliminate_deadend (removing nodes with unused outputs) and eliminate_nop_transpose (removing transpose nodes that do nothing). Graph rewrite passes modify the graph structure to ensure compatibility; for example, rewrite_input_dtype adjusts input data types. In this category, there are also secondary passes designed primarily for analysis and debugging, such as split_init and split_pre-
dict
, which isolate the initializer and inference subgraphs, respectively. By default, the optimizer applies only fusion and elimination passes; however, it also supports applying any subset of its $47$ transformation passes, based on its user-defined configuration.
2.4. Model Output Comparison Methods
Based on the model and output type, OODTE uses a set of well-known and established methods in order to compare the outputs of the optimized model versions with their original counterparts. In particular:
Classification models::
We compare the top-K label predictions between original and optimized models using Kendall’s Tau Rank Correlation (kendall, ), which evaluates the similarity between ranked label lists, with scores ranging from -1 (inverse correlation) to 1 (perfect correlation).
Object Detection & Semantic Segmentation models::
Ea-
ch prediction includes (1) class probabilities, (2) bounding box coordinates, and (3) a confidence score. For class probabilities, we apply a softmax and use Kendall’s Tau to compare rank similarity. For bounding boxes and confidence scores, we use standard metrics: (I) IoU for bounding box overlap, (II) F1 Score for overall detection balance, (III) Average Precision (AP) for precision across thresholds, (IV) Average Recall (AR) for recall across thresholds, (V) mAP for aggregate detection and localization accuracy. These metrics are computed at IoU thresholds of $0.5$, $0.75$, and $0.9$, with the original model as the reference.
Text generation models::
We convert token outputs to plain text and use the BLEU Score (BLEU, ) to compare the optimized model’s output against the original, assessing n-gram overlap and overall similarity.
Binary sentiment analysis models::
We measure the percentage of prediction label differences between original and optimized models to evaluate the impact of optimization on classification consistency.
3. System Architecture & Methodology
Figure 1. Architecture of OODTE  with the following primary modules: (1) Model Orchestrator, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module.
The architecture of OODTE and its modules is shown in Figure 1. Its workflow is straightforward: load models, optimize them, execute both original and optimized versions across the dataset, and compare the results. If output discrepancies are detected, OODTE iteratively re-optimizes each model using individual ONNX Optimizer passes to localize the pass causing the difference. All results are stored in report files for analysis and fault localization.
OODTE consists of four primary modules: (1) Model Orchestrator, (2) Optimizer Module, (3) Runner Module, and (4) Metrics Comparison Module. It loads model configuration settings either (1) from a local configuration file or (2) by fetching them from the ONNX repository from which the models are sourced, both managed by the Model Orchestrator Module. The modules in Figure 1 are detailed in the following sections.
3.1. Model Orchestrator Module
The Model Orchestrator Module is responsible for all tasks related to fetching, loading (including model configuration) and preparing models for optimization and execution. The module supports two loading options for DNN models: (1) local loading of ONNX models, and (2) loading models from an online repository. At the moment, OODTE supports the official ONNX Model Hub (onnxmodelhub, ), however, as it utilizes the official Hub API, it can trivially support third party ONNX model repositories. In addition, the module supports fetching and loading of multiple models simultaneously, while it also performs model caching when fetching from a model hub, as well as checksum verification. Once each model and its respective configuration are properly loaded, they are propagated to the Optimizer Module and the Runner Module.
3.2. Optimizer Module
The Optimizer module is responsible for the generation of optimized models, using the ONNX Optimizer at its heart. It supports two settings: (1) full optimization, applying all the ONNX Optimizer passes to the model under test, and (2) partial optimization, where only a select set of passes is applied to the model, applied in a user-defined order specified in the system configuration. It is worth noting that in (1), by optimizer design, only the fuse and eliminate passes are applied as part of the default optimization bundle in the ONNX Optimizer, leaving out a number of secondary passes (e.g., ‘rename_input_output‘). OODTE supports the utilization of these passes for testing, in order to explore all optimizer operations. This feature is essential throughout the OODTE pipeline: when faulty behavior is discovered during model evaluation, the models exhibiting such faults must be examined in depth. To identify the pass responsible for problematic behavior, OODTE first compares fully optimized models with their originals throughout its pipeline, and upon detecting a crash or accuracy issue, isolates the problematic cases by generating per-pass optimizations and comparing them again with the originals in an iterative process, until the pass responsible for the issue is identified. Once a model is optimized with either setting, it then gets propagated to the Runner model for inference.
3.3. Runner Module
The Runner model does precisely what its name suggests: executes the models under test. In particular, it receives both the original and the optimized model under test, as well as a dataset that will be used as input for both model versions. It then performs model inference, using ONNX Runtime (onnxruntime, ), combined with the necessary model preprocessing, while utilizing the model configuration loaded in the Model Orchestrator component. OODTE supports different types of datasets, based on the type of models being examined - from images to IMDB reviews. It can load the inputs from local files, but also fetch the dataset from well-known packages, such as the datasets (datasets, ) package.
To enable the use of large-scale datasets and extract results from models that are larger and more computationally demanding—especially when operating on low-end devices—the Runner Module supports execution across a dataset in parts. In particular, a run can be divided into $N$ chunks, where each chunk consists of a fraction, $1/N$th, of the dataset. Any remaining data not evenly divisible is also included in the final chunk of the run. This process applies to both the original and the optimized model versions, with the results extracted and propagated to the Metrics Comparison Module for subsequent evaluation and reporting. By following this approach, we were able to conduct our survey on two consumer-grade devices without relying on a compute cluster to obtain results, as described in Section 5.
3.4. Metrics Comparison Module
Once the run completes per-chunk for both the original and the optimized model, the results are propagated to the Metrics Comparison Module. This module performs comparison against the results, by selecting the appropriate comparator based on the model type. For classification types, OODTE supports first-label comparison, but also calculates the Kendall Rank Correlation Coefficient across the topKtop-K predictions - commonly known as Kendall’s Tau (kendall, ). For Object Detection models, it utilizes the same methods of classification for labels and ranks, but also applies Intersection over Union (IoU (IoUBoundingBox, )), Precision, Recall, and F1 Score (F1Score, ) given an IoU threshold. For text generation models (e.g., GPT-2), it utilizes BLEU (BLEU, ) score in order to compare the text generated out both the original and the optimized model, as well as direct tensor comparison for cases of question answering (Q&A). In addition OODTE is build in a manner that new comparators can be added, in cases that a new category of models is tested.
3.5. Fault Localization
Once a method is selected, the Metrics Comparison Module calculates the metrics based on model type and generates a report with all results, in JSON format. If a difference is detected across the original and the optimized model, then OODTE retrofits the reported data to the Optimizer Module. Then, the Optimizer Module generates model optimizations per-pass, reporting how each pass behaves. OODTE applies a fault localization methodology for three cases: when the optimizer crashes, when it generates an optimized model in an invalid form, and when the optimization succeeds but produces results inconsistent with the original model. In addition, OODTE registers any warnings that can indicate the presence of a potential performance fault (e.g., redundant initializers in the model). Upon detecting an issue or a problematic behavior, OODTE attempts to identify the responsible pass by generating per-pass optimizations, executing inference on the resulting models, and comparing their outputs with those of the original model. This process is applied iteratively for each of the passes supported by the optimizer, until the pass responsible for the detected issue is identified.
4. Related Work
In this work, we compare our work with contributions related to AI compilers, focusing on the following key aspects: (1) the ability to localize faults associated with graph optimization passes, (2) the detection of compiler bugs affecting The types of bugs considered are the ones resulting in (a) model accuracy (inference results), (b) compiler crashes, (c) successful compilations, which, however, produce malformed or incorrect models that fail to run (e.g., models with incorrect input nodes or parameters), and (d) issues impacting execution times, (3) the capability to localize faults in the ONNX domain, particularly within ONNX Optimizer, and (4) the detection of faults in graph optimization passes. Table 1 summarizes the features supported by each contribution. For each tool, Table 1 shows: (1) the underlying technique, (2) support for bug detection in ONNX, (3) the ability to localize faults (Flt Loc.), (4) the types of bugs detected and the corresponding graph optimization passes involved (Gr. Opt. Passes), (5) detection of label accuracy faults (Acc.), (6) detection of crash faults (Cr.), (7) detection of malformed models produced despite successful compilation (Malf.), (8) detection of bugs affecting execution times (Ex. T.), and (9) the compilers supported/evaluated.
Table 1. Comparison of features between OODTE and related work. Each technique is compared across the following parameters: testing of ONNX (ONNX?), ability for fault localization (Flt. Loc.), bug types, graph optimization passes (Gr. Opt. Passes), detection of accuracy changes (Acc.), crashes (Cr.), generation of malformed models due to miscompilation (Malf.), and issues affecting model execution times(T.) bugs, as well as the compilers explored on each contribution.
:*: Limited support of feature, $\dagger:$ Does not consider ONNX Optimizer or compile-time optimizations associated with ONNX, $\ddagger$}: Uses ONNX as an intermediate representation, but does not perform fault detection associated to it.
Tool Technique ON NX? Flt. Loc. Bug Types Gr. Opt. Passes Acc. Cr. Malf. T. Compilers
OODTE
Per-pass fault
localization,
differential
testing
Logic, API versioning,
graph input/parameters/
transformations
Fuse/
eliminate/
rewrite
ONNX
Optimizer (onnxoptimizer, )
NNSmith (NNSmith, )
Grammar-
based fuzzing
†,‡ $\times$
Conversions, semantics,
gen.transformations,
graph rewrites
Conv2D
inputs
rewrite
$\times$ $\times$
TVM (tvm, ),
PyTorch JIT (pytorch, ),
TensorRT (tensorrt, ),
PT2 (pytorch2, ), ONNX
Runtime (onnxruntime, ),
TF XLA (tfxla, ),
TFLite (tensorflow, )
MLIRSmith (MLIRSmith, )
Grammar-
based fuzzing
$\times$ $\times$
Conversions (dialects),
generic transformations,
logic (dialects)
$\times$ $\times$ $\times$ $\times$
MLIR (MLIR, ),
LLVM (LLVM, )
MLIRod (SuoMLIR, )
Operator
dependency
analysis/
fuzzing
$\times$ $\times$
Conversions (dialects),
gen. transformations
$\times$ $\times$ $\times$ $\times$ MLIR (MLIR, )
SYNTH-
FUZZ (synthfuzz, )
Grammar-
based fuzzing
$\times$ $\times$ Generic, input-based $\times$ $\times$ $\times$ $\times$ MLIR (MLIR, )
HirGen (HirGen, )
Graph IR
fuzzing
$\times$ $\times$
Numeric computations,
exceptions, types, memory
$\times$ $\times$ $\times$ $\times$ TVM (tvm, )
TVMFuzz (TVMFuzz, )
Low-level
IR fuzzing
$\times$ $\times$
Numeric computations,
graph rewrites
Constant
folding
$\times$ $\times$ TVM (tvm, )
Tzer (TZER, )
Coverage-
guided fuzzing
(Low-level IR)
$\times$ $\times$
API inconsistency/misuse,
memory, types
$\times$ $\times$ TVM (tvm, )
GenCoG (GenCoG, )
Constraint
solving,
High-level
IR Fuzzing
$\times$ $\times$
Memory, tensor shape
/data, operator attributes,
logic
Operator
fusion
$\times$ $\times$ TVM (tvm, )
DeepDiffer (DeepDiffer, )
Low-level IR/
coverage-
guided fuzzing,
differential
testing
$\times$ $\times$
Type, memory, API misuse,
numeric computations,
exception handling
$\times$ $\times$ $\times$ TVM (tvm, )
WhiteFox (WhiteFox, )
LM analysis,
fuzzing
$\times$ $\times$
Memory, API misuse,
concurrency, logic,
security vulnerabilities
$\times$ $\times$
TF XLA (tfxla, ),
PyTorch
Inductor (pytorch, ),
LLVM (LLVM, )
TorchProbe (Su_2023, )
Feature-
based fuzzing
$\times$ $\times$
Memory, API misuse,
logic
$\times$ $\times$ $\times$
PyTorch (pytorch, ),
Triton (TilletTriton, )
MT-
DLComp (Xiao_2022, )
Metamorphic
testing
$\times$ $\times$
Memory, dead code,
layout optimization,
numeric accuracy
Conv2D
layout
change
$\times$ $\times$
TVM (tvm, ),
Glow (Glow, ),
NNFusion (NNFusion, ),
TF XLA (tfxla, )
FetaFix (FetaFix, )
Fault localiza-
tion,
differential
testing
$\times$
Preprocessing, input
dimensions, parameters,
hyperparameters, tensor
shape, graph structure
$\times$ $\times$ $\times$ $\times$
Converters across
PyTorch (pytorch, ),
Keras (chollet2015keras, ),
TF (tensorflow, ), TFLite (tensorflow, )
In terms of the Technique column shown in Table 1, most existing works primarily rely on compiler fuzzing, often enhanced with techniques such as differential testing (e.g., DeepDiffer (DeepDiffer, )) or LLM analysis (e.g., WhiteFox (WhiteFox, )). Consequently, these systems focus on triggering compiler bugs rather than pinpointing their origins and root causes, typically relying on generic compiler error messages to perform manual fault localization. Among these works, only FetaFix (FetaFix, ) provides a mechanism for fault localization, but it targets accuracy-related faults in model conversion between DL frameworks and does not support the model optimization process. In contrast, OODTE systematically localizes faults by applying an iterative fault localization methodology for each pass in the ONNX Optimizer.
In terms of bug types detected in Table 1, existing tools primarily focus on detecting crash bugs. For label accuracy-related bugs, they typically rely on random input generation (e.g., NNSmith (NNSmith, )) to expose inconsistencies in model inference. In contrast, for label inconsistency bugs that OODTE detects, we assess the prevalence and impact of these bugs on popular datasets(e.g., ILSVRC (ILSVRC17, ), IMDB (imdbsentiment, )) and model architectures. Additionally, only a small number of tools consider miscompilations (for instance, models that contain problematic input nodes or misconfigured parameters), with just two prior works capable of detecting faults of this kind (TVMFuzz, ; WhiteFox, ). The detection of bugs affecting execution times of the compiled models has received limited attention in the literature, with only two TVM-focused systems (TZER, ; DeepDiffer, ) using empirical measurements to evaluate performance-related faults. OODTE currently identifies issues that could potentially lead to degraded model execution times by flagging redundant operator warnings (e.g., unused initializer entries), while the integration of precise execution time measurements is left as subject for future work.
Regarding support for ONNX in Table 1, OODTE is the only tool capable of localizing faults in the ONNX Optimizer. Other methodologies either treat ONNX merely as an intermediate representation or focus on bug detection in ONNX Runtime (WhiteFox, ), a cross-platform inference engine, unlike model optimization targeted by the ONNX Optimizer.
Existing work on bug detection within graph optimization passes (shown as Gr. Opt. Passes column in Table 1) is limited, with prior approaches focusing on common operations such as constant folding, operator fusion, and convolution optimizations. In contrast, OODTE addresses the full spectrum of features in the ONNX Optimizer, covering a set of 47 passes that perform a wide variety of graph optimizations such as Fuse, Eliminate, and Rewrite operations in addition to the more common operations.
5. Experiment Setup
5.1. The Experiment Process
We evaluated all ONNX Optimizer passes (Section 2.3). For each model, we first applied the default setting, which runs all fusion and elimination passes, and compared the outputs of the original and optimized models as described in Section 3.4. When discrepancies arose, we re-optimized using individual passes, producing 47 single-pass models to identify the pass causing the issue. OODTE also supports new or custom passes, and our methodology generalizes to other optimization frameworks (e.g., Apache TVM) that allow both bundled and individual pass execution.
5.2. DNN Models
For our experiment set, we sourced $130$ model instances from the ONNX Model Hub (onnxmodelhub, ). In particular, we selected $93$ models utilized for classification tasks (e.g., MobileNet (mobilenet, ), EfficientNet-Lite4 (efficientnet, ) DenseNet121 (densenet, ), and InceptionV2 (inception, )), $30$ models used for object detection and semantic segmentation tasks (e.g., RetinaNet (retinanet, ), Mask/Faster R-CNN R-50-FPN (mask-rcnn, ; faster-rcnn, ), and SSD (ssd, )), and $7$ transformer models (attention, ) utilized for text comprehension and generation tasks, such as question answering (Q&A), (BERT-SQuAD (BERT, )), sentiment analysis (RoBERTa (liu2020roberta, )), text summarization (T5 (T5Paper, )) and generic-purpose text generation (GPT-2 (gpt2, )), with $4$ models focusing on Q&A, while one model is utilized for each of the other text comprehension and generation tasks. We utilized models of a variety of opsets based on hub availability, ranging from $7$ (the lowest compatible with the latest ONNX Runtime version by the time of this publication - v1.21.11.21.1) up to $12$ (the highest provided from ONNX Model Hub via their API). Our methodology, however, is applicable to any ONNX model type and is not constrained by opset versions or tied to a specific model repository.
5.3. Computational Resources
We ran our models on mid-range systems with Intel i5 and i7 CPUs, reflecting the typical hardware used by developers during model development and fine-tuning. Our method is generalizable to larger models, limited only by ONNX model size constraints, and can be extended to high-end systems with a trivial extension for hardware acceleration (using the ONNX Runtime API (onnxruntime, ), or an AI compiler such as Apache TVM (tvm, )).
5.4. Datasets
We selected datasets appropriate to each model task. For classification and object detection, we used the ILSVRC2017 (ILSVRC17, ) validation dataset, containing $5500$ images, commonly used for benchmarking image recognition models (e.g., ResNet (resnet, ) and BERT (BERT, )). For text comprehension and generation, we used SQuAD-v2 (rajpurkar2018knowdontknowunanswerable, ), which includes $11873$ questions and contexts, to evaluate Q&A tasks and general-purpose transformer behavior. Additionally, for summarization and sentiment analysis, we used the IMDB dataset (moviedataset, ; imdbsentiment, ), containing $25000$ entries.
To accurately detect differences across model types, we employed comparison methods appropriate for each task. Specifically, we used the following approaches:
  1. (1)
    For image recognition models performing classification tasks, we utilized the Kendall Rank Correlation Coefficient (kendall, ) (KT), measuring differences for top-K prediction labels of each dataset input image, for $K=1,5$, and $10$. If the KT value was found to be <1<1 for each input, we counted this input as different.
  2. (2)
    For image recognition models performing object detection and semantic segmentation, we calculated precision, recall, and F1 score for each dataset input, and extracted their average, combined with different thresholds of Intersection-over-Union (IoU) ($IoU=0.5,0.75$, and $0.9$), as well as IoU, mean Average Precision (mAP), along with average precision and recall. We considered the original model output as the base, and the optimizer model the one under test for our metric calculations, and the higher the values observed, the better the optimized model aligns with the original.
  3. (3)
    For text generation models performing question answering and summarization, we utilized BLEU score (BLEU, ) across output texts for each input, monitoring instances where BLEU_Score<1BLEU\_Score<1 for each input.
  4. (4)
    For sentiment analysis models, we compared each output value consisting of 0,1{0,1} across the original and the optimized versions, registering any differences observed.
6. Results
Following the execution of the full set of experiments. A complete list of these issues, along with the associated models, is shown in Table 2. The developers have acknowledged the issues we reported, and we have provided them with full reproducibility details for each reported bug. Of the $15$ bugs identified, $14$ were entirely new (not previously reported on the ONNX Optimizer issue tracker), while one was a previously reported issue that lacked sufficient detail, including model information and reproducibility steps. We provided comprehensive information and full reproducibility details for this issue.
Within the ONNX optimizer, the bugs identified were primarily associated with $9$ of the $47$ passes. However, not all passes are equally important. By default, the optimizer prioritizes passes related to fuse and eliminate operations. As a result, $8$ of the $11$ reported GitHub issues are considered more critical since they belong to the optimizer’s core (default) strategy. Among these $8$ issues, $6$ are linked to specific passes (e.g., fuse_bn_into_conv, which fuses batch normalization nodes with convolutions), while the remaining 2 issues were more general and related to the optimizer’s overall usage. We analyze and present the impact of these faults both on the optimization process and on the correctness of the optimized models in the following Sections.
6.1. Optimizer Crashes
Out of the $15$ detected bugs, $12$ bugs were related to model crashes, $2$ bugs resulted in output differences between the original and the optimized model, and $1$ bug introduced redundant initializers in the model that produced runtime warnings. In particular, we observed crashes were caused by $8$ out of the $47$ passes (1717%), and occurred for specific models, with object detection models being more prone to errors from these $8$ passes. In particular, R-CNN models, such as Faster R-CNN R-50-FPN (faster-rcnn, ), RetinaNet (retinanet, ), VGG-16, VGG-19 (vgg, ), as well as YOLOv3 (Redmon2018YOLOv3AI, ), presented crashes by using the optimizer, in a range of opset instances and in passes related to the modification of gather nodes, and the elimination of idempotent operations. In total, we observed $31$ model instances where either the optimizer crashed upon processing them, or the optimized model was problematic, leading to a runtime crash. Some secondary passes are known to be unstable (e.g., split_init and split_predict describe in their documentation that they not strictly comply with the ONNX format), and therefore we do not focus on them as sources of faults in the optimization process. Overall, $12$ out of the $31$ model instances (9.29.2%) that were associated with a crash occurred while using passes related to fuse and eliminate operations, which are part of the default bundle of optimizations in the ONNX Optimizer and are therefore of particular significance.
As shown in Table 2, crash faults can be attributed to three main causes: (1) model input issues (cases #$4$ and #$6$), (2) incorrect use of initializers (cases #$12$ and #$13$), and (3) issues associated with node properties such as duplicate names, obsolete fields, and incorrect dimensions (all other cases). Overall, object detection models were more prone to optimizer crashes compared to classification models. Due to their increased structural complexity and larger size, they are more likely to exhibit errors related to node properties and initializers, which account for the vast majority of crash faults. In contrast, only one issue was observed in the most complex model category — text generation models, that was related to model input in GPT-2. However, upon further inspection, we found that the transformation passes of the ONNX Optimizer had limited impact on text generation models — the majority of optimization passes left the models unchanged, and those that did make modifications affected a small proportion of nodes with negligible effect on outputs.
Table 2. Overview of the issues identified using OODTE, detailing the responsible optimization pass, the detected error, its type (Crash, Accuracy (Acc.), or Warning (Warn.)), and the models in which the error was observed, along with their opsets (op.).
# Pass Error Type Models
1 lift_lexical_references
Removes used
value references.
Crash
EfficientNet-Lite4 (op. 11), SSD (op. 12)
2 split_init
List index
out of range.
Crash YOLOv3 (op. 10, 12)
3
split_init
split_predict
Required Field
is missing.
Crash
DenseNet121 (op. 7-9), YOLOv2 (op. 10),
SSD (op. 12), Tiny-YOLOv3 (op. 11),
YOLOv3 (op. 10, 12)
4 split_predict
Required inputs
are missing.
Crash
SSD (op. 12), DenseNet121 (op. 7),
InceptionV2 (op. 6-9)
5 split_predict Bad spec for node. Crash
YOLOv3 (op. 10), DenseNet121 (op. 7),
InceptionV2 (op. 6-9)
6 split_predict
Mismatch in number
of input elements
Crash
FCN ResNet-50-int8 (op. 12)
7 eliminate_shape_gather
Indices value
not equal to
node dims size
Crash
Faster R-CNN: R-50-FPN (op. 10)
8 All Passes
Implicit ONNX
format version
update.
Acc.
Models of cases 11-$11$
in Table 3
9 rewrite_input_dtype
Output name
used multiple times.
Crash GPT-2 (op. 10)
10
eliminate_consecutive_idempotent_ops
Incompatible shape
error.
Crash
Mask R-CNN/Faster R-CNN: R-50-FPN-int8 (12),
R-50-FPN-qdq (op. 12), R-50-FPN-fp32 (op. 12)
11
eliminate_duplicate_initializer
Erase output
assertion failed.
Crash
Faster R-CNN: R-50-FPN-qdq (12)
12 fuse_bn_into_conv
Duplicate initializer
introduced.
Crash
RetinaNet (op. 9)
13
eliminate_shape_gather
eliminate_nop_reshape
Added initializers
not used by the model.
Warn.
Faster R-CNN: R-50-FPN-int8 (op. 12)
14 All Passes Invalid input scale. Crash
VGG-16-BN (op. 7), VGG-19-BN (op. 7),
ResNet101_DUC_HDC (op. 7)
15 fuse_bn_into_conv Bad graph structure. Acc.
Efficient-Lite4 (op. 11),
SSD (op. 12)
6.2. Output Label Differences in Optimized Classification & Object Detection Models
Out of the $130$ optimized models tested, we observed unexpected output differences between the original and optimized models. These differences manifested as (1) label differences for classification models (e.g., ResNet), (2) label and/or bounding box differences for object detection models (e.g., YOLOv3), and (3) textual output differences in token sequences or predicted spans for text generation tasks. Since the ONNX Optimizer transformations are graph-based and do not alter numeric precision, output differences are not expected. Among the $93$ classification models tested, $28$ (30.130.1%) exhibited discrepancies in output labels. Similarly, $5$ out of $30$ (16.616.6%) object detection models presented output differences. The text generation models, on the other hand, were robust with only one instance (GPT-2) exhibiting negligible differences ($0.001\%$) while all others showing identical results across the original and the optimized model versions. The differences impacted the top-5 and top-10 predictions, while top-1 predictions remained unchanged. Although top-1 accuracy is often emphasized, variations in lower-ranked predictions can also reveal important issues. Our results show that focusing exclusively on top-1 accuracy may give developers a false sense that an optimization is successful, while overlooking substantial changes in top-K predictions (up to 2% in top-5 and 13% in top-10). Using OODTE, we detected these inconsistencies and reported them to the ONNX Optimizer developers.
Table 3. Output label accuracy differences across classification (CL) and object detection (OD) models, showing type, opset version (Ver.), and accuracy gaps for top-1 (T-1), top-5 (T-5), and top-10 (T-10) predictions between original and ONNX-optimized models, along with the related issue (Is.): Format Version (FV) or Malformed Graph (MG).
# Model Type Ver. T-1 T-5 T-10 Is.
1 AlexNet (alexnet, ) CL 7-9 0% 0% 0.02% FV
2 DenseNet121 (densenet, ) CL 7-9 0% 0% 0.02% FV
3 GoogleNet (googlenet, ) CL 7-9 0% 0% 0.04% FV
4 Inception-1 (inception, ) CL 7-9 0% 0% 0.25% FV
5 Inception-2 (inception, ) CL 7-9 0% 0.02% 0.05% FV
6 MobileNet v2 (mobilenet, ; mobilenetv2, ) CL 7-9 0% 0% 0.02% FV
7 ResNet18 (resnet, ) CL 7 0% 0.02% 0.02% FV
8 ResNet34-v2 (resnetv2, ) CL 7 0% 0.02% 0.02% FV
9 ResNet50-caffe2 (resnet, ) CL 7-9 0% 2% 13% FV
10 ShuffleNet-v1 (shufflenet, ) CL 7-9 0% 0.7% 1% FV
11 SqueezeNet 1.0 (squeezenet, ) CL 7-9 0% 0% 0.02% FV
12 EfficientNet-Lite4 (efficientnet, ) CL 11 0% 0.04% 1% MG
13 SSD (ssd, ) OD 12 0% 0% 0.04% MG
Overall, we detected two primary issues related to output label differences between the original and optimized models:
Implicit ONNX format version upgrade::
The ONNX Optimizer silently upgrades models from ONNX format version v3 to v4, without any warning or indication. This behavior impacts models using opset versions 7, 8, and 9— that are older opsets but still supported in the current ONNX Runtime (v1.21.1). Despite being outdated, these opset versions are still prevalent, representing approximately 4141% of the models in the ONNX Model Hub (presented in Table 3, with the issue labeled as “FV”, indicating a problem with the format version.). While not all models within this opset range are affected, we observed a significant number that exhibited differences between their original and optimized versions. We have reported this issue to the ONNX Optimizer developers, who have acknowledged it and are investigating further.
Problematic fusion pass::
The pass fuse_bn_into_conv
fuses Batch Normalization and Convolution operators within the model. In particular, the pass attempts to fuse BatchNorma-
lization
nodes (which do not update their mean or variance during inference) with their preceding Conv nodes by adjusting the weights and biases of the Conv node according to the properties of the BatchNormalization node. We observed that this pass was problematic in a number of cases, resulting in both a crashing bug and an issue where redundant nodes are introduced into the model. In the latter scenario, instead of properly merging the nodes resulting in a Conv node with adjusted weights and biases, the pass generates multiple nodes that represent the Batch Normalization formula. This behavior is illustrated on the right side of Figure 2. This formula representation is problematic as several models produced different output labels after optimization using this pass compared to their original versions (presented in Table 3, with the issue labeled as “MG”, indicating a malformed graph). Although another user previously reported a similar issue, no details on the reproduction steps or affected models were provided. We submitted a reproducible issue to the ONNX Optimizer developers, including all relevant information and behavior details.
Table 3 shows label differences in top-K outputs for classification and object detection models, highlighting the causes.
Table 4. Object detection metric differences between the original and optimized versions, using the default ONNX Optimizer setting. For each model, we present the average F1 score (F1), precision (Prec.), and recall (Rec.), as well as mAP and IoU.
Model Op. F1 Prec. Rec. mAP IoU
SSD (ssd, ) 12 99.99% 99.99% 99.99% 99.99% 100%
Tiny
YOLOv3 (Redmon2018YOLOv3AI, )
11 99.18% 99.18% 99.18% 99.99% 89.35%
YOLOv2 (yolov2, ) 9 99.95% 99.95% 99.95% 99.94% 88.49%
YOLOv3 (Redmon2018YOLOv3AI, ) 10 99.45% 99.45% 99.45% 99.99% 84.81%
YOLOv3 (Redmon2018YOLOv3AI, ) 12 99.45% 99.45% 99.45% 99.99% 84.81%
Figure 2. Incorrect model graph transformation caused by the fuse_bn_to_conv pass in EfficientNet-Lite-4. The nodes affected are highlighted. It should produce a Conv node with updated weights and biases, but the pass generates multiple nodes that represent the BatchNormalization formula.
Bounding Box Differences in Object Detection Models We observed some differences in the bounding box outputs between the original and optimized versions of object detection models. The results are presented in detail in Table 4 with the following main findings: 1. Both YOLOv3 models (opset $10$ and $12$) achieve high average IoU, F1 score, precision, and recall, all above 99.4599.45%. However, the optimized model occasionally misplaces bounding boxes—resulting in IoU drops of up to 15.1915.19%, these differences do not lead to misclassifications, as none were observed in the top ranked label predictions. A similar trend is also seen in YOLOv2 model where optimization causes a drop in IoU value but not in detection performance (F1, precision and recall).
6.3. Optimizer Warnings
We observed two issues related to the addition of redundant elements (unused initializer entries), introduced in the optimized models, associated with the eliminate_shape_gather and eliminate_nop_reshape passes. When executed with ONNX Runtime, these entries were ignored, triggering a warning in the format: Removing initializer ‘INITIALIZER_ID’. It is not used by any node and should be removed from the model. Upon manual inspection, we found no significant differences in model size. However, this issue warrants further investigation for potential side effects towards model performance. We have reported it to the optimizer developers for additional review.
7. Tool Usability
OODTE  can test whether the ONNX Optimizer generates optimized models without compromising output label correctness. Currently, OODTE supports differential testing for classification, object detection, and text generation models. However, OODTE can also be easily extended to support new model types, as long as (1) the models can be compiled into the ONNX format for optimization and (2) a comparator is provided to compare the results between the base and optimized models for the new model type. In addition, OODTE can fetch models and datasets automatically, making it adaptable for use in a Continuous Integration pipeline. This allows OODTE to test the ONNX Optimizer whenever new functionality (e.g., new passes) is added. By default, OODTE loads models from the official ONNX Model Hub (onnxmodelhub, ), which was used in our experiments. However, by configuring OODTE to use a custom ONNX Hub that follows the ONNX API specification for model repositories (onnxhubapi, ), it can fetch and use models without further modification. The same flexibility applies to datasets. OODTE can fetch datasets using packages like the datasets (datasets, ) package or from local storage.
Finally, while OODTE applies a differential testing methodology focused on the ONNX Optimizer, this approach is generalizable and can be easily adapted to other AI optimizers or compilers that meet two primary requirements: (1) they provide a runtime for model execution, and (2) they expose an API that includes all the passes applied during model optimization, both as a bundle and as independent passes.
8. Threats To Validity
We list the threats to validity in our evaluation below:
Model Selection: Our evaluation covered 130 models from the official ONNX Model Hub, spanning classification, object detection, as well as text comprehension and generation. Models were chosen for their maturity, accessibility, and broad adoption. aimed to create a representative and diverse experimental set, although the selection is not exhaustive. We also excluded models with opset versions below 7 due to incompatibility with current ONNX and ONNX Runtime versions, limiting our results to newer models. To keep the study tractable, diffusion and reinforcement learning models were omitted; we plan to include these in future work.
Methodology Generalizability: We focus ONNX Optimizer because it is a core component of the widely used ONNX ecosystem (19.5K stars) (onnx, ) and supports tasks beyond model optimization, such as graph simplification (ONNX Simplifier, 4.1K stars (onnxsimplifier, )). Its popularity and broad applicability make it a strong candidate for study and a representative of the state-of-the-art technologies. While testing all DL compilers was infeasible, we demonstrated the value of our methodology and noted it can extend to other optimizers that offer per-pass APIs and runtime support for differential testing. For example, OODTE could support OpenVINO (openvino, ) and Apache TVM (tvm, ) by adapting their graph optimization passes, considering both compilers also provide runtime APIs, similarly to ONNX.
Optimization Pass Ordering:
The order of optimization passes can significantly affect model performance and behavior (OptimizationOrder, ), a well-known challenge in compiler research. While we do not study pass ordering in depth, OODTE adopts two evaluation modes: (1) the default ONNX Optimizer configuration, including common passes such as fuse and eliminate, and (2) a configuration applying all 47 available passes atomically. Although this does not capture issues from specific pass combinations or orderings, it effectively uncovers correctness problems in practice. Nevertheless, the omission of a systematic study on optimization ordering may limit the completeness of our evaluation.
9. Discussion
We summarize below our findings for the different model types in our experiments.
(1) Optimizations can affect label accuracy in classification models. We identified two primary causes for label accuracy differences after optimization: (a) Legacy opset upgrade: Some models used older ONNX opsets, and the optimizer implicitly and improperly upgraded their format versions. Classification models built on these legacy opsets were particularly affected. (b) Problematic graph node generation: The fuse_bn_into_conv pass, which fuses batch normalization into convolutional layers, can unintentionally alter model behavior and particularly affects classification models that rely on convolutions.
Although text generation and comprehension models are also large, manual inspection showed that the optimizer applies only a few passes to them. Most optimizations are skipped because the targeted layer patterns (e.g., convolutions) are uncommon in transformer-based architectures used for language tasks.
(2) Object detection models crash more frequently when optimized.
Most crashes during optimization occurred in object detection and semantic segmentation models, rather than in classification models. This is likely due to their greater architectural complexity—being larger and structurally more complex—which leads the ONNX Optimizer to apply more extensive transformations, increasing the likelihood of encountering corner cases that trigger crashes related to the ONNX Optimizer passes.
(3) Optimization can affect bounding boxes of object detection models. Regarding accuracy, object detection models remained largely stable. We observed only a negligible decrease in the accuracy of SSD on top-10 accuracy predictions (0.040.04%), while the other models showed no difference between the original and optimized versions. However, the optimized models occasionally misaligned bounding boxes, leading to drops in IoU value of up to 15.1915.19%.
(4) Outputs from text comprehension and generation models were not affected by optimization. Among the 8 text comprehension and generation models tested, we observed only one crash (in GPT-2) and a small number of samples (8) with BLEU-score differences between the original and optimized models for the top-3 token selections.
Overall, we find ONNX optimizations can result in output differences between the optimized and unoptimized model, with some model types being more susceptible to optimization errors than others. OODTE was able to successfully isolate the optimization passes responsible for the output differences for different model types.
10. Conclusion
We propose OODTE, a tool for the differential testing of the ONNX Optimizer. Using OODTE, we evaluated $130$ models across classification, object detection, and text generation tasks. We detected 9.29.2% of model instances that crashed when using the primary optimizer strategies, while 3030% of classification models and 16.616.6% of object detection models showed differences between the source and optimized versions, while text generation models remained robust. Additionally, OODTE identified 15 issues, 14 of them new, related to crashes and accuracy degradation across 9 optimizer passes, which we reported to the ONNX Optimizer developers.